### v2

# A logifold structure of measure spaces

Abstract.
In this paper, we develop a local-to-global and measure-theoretical approach to understand datasets. The idea is to take network models with restricted domains as local charts of datasets. We develop the mathematical foundations for these structures, and show implementations how it can be used to find fuzzy domains and to improve accuracy in data classification problems.
1. Introduction
Let $X$ be a topological space, $B_{X}$ the corresponding Borel $\sigma$-algebra, and $\mu$ a measure on $(X,B_{X})$. In geometry and topology, the manifold approach dated back to Riemann uses open subsets in $\mathbb{R}^{n}$ as local models to build a space. Such a local-to-global principle is central to geometry and has achieved extremely exciting breakthroughs in modeling spacetime by Einstein’s theory of relativity.
In recent years, the rapid development of data science brings immense interest to datasets that are ‘wilder’ than typical spaces that are well-studied in geometry and topology. Moreover, the development of quantum physics also raises a challenging question of whether the spacetime can be treated as a manifold.
Taming the wild is a central theme in the development of mathematics. Advances in computational tools have helped to expand the realm of mathematics in the history. For instance, it took hundreds or thousands of years in human history to recognize irrational numbers and to approximate $\pi$ by rational numbers. From this perspective, we consider machine learning by network models as a modern tool to express a ‘wild space’ as a union (limit) of spaces that admit a finite expression.
From a perspective of small scale, a dataset can be thought as a union of numerous disperse tiny open balls (‘thickened points’). We can regard such a union as a manifold with many contractible components. However, this is not so useful in applications since this forgets about the relations between the different components. In a larger scale, a dataset is a measure space in nature. We would like to formulate local charts for a measure space that capture more information and admit finite mathematical expressions. Such an atlas of local charts endows a measure space with an additional geometric structure.
In this paper, in place of open subsets of $\mathbb{R}^{n}$, we formulate ‘local charts’ that are closely related to network models and have logic gate interpretations on (fuzzy) measure spaces. Network models provide a surprisingly successful tool to find mathematical expressions that approximate a dataset. Non-differentiable or even discontinuous functions analogous to logic gate operations are frequently used in network models. It provides an important class of non-smooth and even discontinuous functions to study a space.
In our formulation, the measure of each chart is required to be non-zero. To avoid triviality and requiring too many charts, we may further fix $\epsilon>0$ and require that $\mu(X-U)<\epsilon$. Such a condition disallows $U$ to be too simple, such as a tiny ball around a data point in $X$. This resembles the Zariski-open condition in algebraic geometry.
As a first step, we take classification problems as the main motivation. For this, we consider the graph of a function $f:D\to T$, where $D$ is a measurable subset in $\mathbb{R}^{n}$ (with the standard Lebesgue measure), and $T$ is a finite set (with the discrete topology). The graph $\mathrm{gr}(f)\subset D\times T$ is equipped with the push-forward measure by $D\to\mathrm{gr}(f)$.
We will use the graphs of linear logical functions $f:D\to T$ explained below as local models. Then a chart is of the form $(U,\Phi)$, where $U\subset X$ is a measurable subset which satisfies $\mu(X-U)<\epsilon$, and $\Phi:U\to\mathrm{gr}(f)$ is a measure-preserving homeomorphism. We define a linear logifold to be a pair $(X,\mathcal{U})$ where $\mathcal{U}$ is a collection of charts $(U_{i},\Phi_{i})$ such that $\mu(X-\bigcup_{i}U_{i})=0$.
Linear logical functions are motivated from network models and have a logic gate interpretation. A network model consists of a directed graph, whose arrows are equipped with linear functions and vertices are equipped with activation functions, which are typically ReLu or sigmoid functions in middle layers, and are sigmoid or softmax functions in the last layer. Note that sigmoid and softmax functions are smoothings of the discrete-valued step function and the index-max function respectively. The smoothing can be understood as encoding the fuzziness of data.
For the moment, let’s get rid of fuzziness by replacing the sigmoid and softmax functions by step and index-max functions respectively. Then we can build a corresponding linear logical graph for the network. At each node of the graph, there is a system of linear inequalities (on the input Euclidean domain $D\subset\mathbb{R}^{n}$) that produces Boolean outcomes that determine the next node to go for the signal. We call the resulting function to be a linear logical function.
We prove that linear logical functions can approximate any given measurable function $f:D\to T$, where $D$ is a measurable subset of $\mathbb{R}^{n}$ with $\mu(D)<\infty$. This provides a theoretical basis of using these functions in modeling data.
Theorem 1.1 (Universal approximation theorem by linear logical functions).
Let $f:D\to T$ be a measurable function whose domain $D\subset\mathbb{R}^{n}$ is of finite Lebesgue measure, and suppose that its target set $T$ is finite. For any $\epsilon>0$, there exists a linear logical function $L$ and a measurable set $E\subset\mathbb{R}^{n}$ of the Lebesgue measure less than $\epsilon$ such that $L|_{D-E}\equiv f|_{D-E}$.
By taking the limit $\epsilon\to 0$, the above theorem finds a linear logifold structure on the graph of a measurable function. In reality, $\epsilon$ reflects the error of a network in modeling a dataset.
It turns out that for $D=\mathbb{R}^{n}$, linear logical functions $\mathbb{R}^{n}\to T$ (where $T$ is identified with a finite subset of $\mathbb{R}$) are equivalent to semilinear functions $\mathbb{R}^{n}\to T$, whose graphs are semilinear sets defined by linear equations and inequalities [31]. Semilinear sets provide the simplest class of definable sets of so-called o-minimal structures [31], which are closely related to model theory in mathematical logic.
O-minimal structures made an axiomatic development of Grothendieck’s idea of finding tame spaces that exclude wild topology. On one hand, definable sets have a finite expression which is crucial to the predictive power and interpretability in applications. On the other hand, the setup of measurable sets provides a big flexibility for modeling data.
Compared to more traditional approximation methods, there are reasons why linear logical functions are preferred in many situations. First, when the problem is discrete in nature, it is simple and natural to take the most basic kinds of discrete-valued functions as building blocks. Second, such functions have a network nature which is supported by current computational technology. Third, we can make fuzzy and quantum analogs which have rich meanings in mathematics and physics. In particular, in Section 2.6, we find how averaging in a quantum-classical system leads to non-linear propagation of states.
Figure 1. An example of a logifold. The graph jumps over values 00 and $1$ infinitely in left-approaching to the point marked by a star (and the length of each interval is halved). This is covered by infinitely many charts of linear logical functions, each of which has only finitely many jumps.
Besides discontinuities, another important element is fuzziness. A fuzzy space is a topological measure space $X$ together with a continuous measurable function $\mathcal{P}:X\to[0,1]$ that encodes the probability of whether a given point belongs to the set. In practice, there is always an ambiguity in determining whether a point belongs to a dataset.
A linear logical function is represented by a directed graph $G$. By equipping each vertex of $G$ a state space and each arrow a continuous map between the state spaces, fuzzy linear logical functions (Definition 2.10) and fuzzy linear logifolds (Definition 3.13) can be naturally defined. In this formulation, the walk on the graph (determined by inequalities) depends on the fuzzy/quantum propagation of the internal state spaces.
Figure 2. The left hand side shows a simple example of a logifold. It is the graph of the step function $[-1,1]\to\{0,1\}$. The figure in the middle shows a fuzzy deformation of it, which is a fuzzy subset in $[-1,1]\times\{0,1\}$. The right hand side shows the graph of probability distribution of quantum observation, which consists of the maps $\frac{|z_{0}|^{2}}{|z_{0}|^{2}+|z_{1}|^{2}}$ and $\frac{|z_{1}|^{2}}{|z_{0}|^{2}+|z_{1}|^{2}}$ from the state space $\mathbb{P}^{1}$ to $\{0,1\}$.
Logifold structure for dataset provides geometric language for Ensemble theory. Ensemble theory in decision making utilizes multiple classifiers (also called as hypotheses, or trained moels) to make a decision, see for instance [9]. Ensemble machine learning achieves improvement in classification problems such as [2] and [3]. In section 4, we briefly introduce existing methods and present a unified approach to Ensemble theory using the logifold structure with the theorem 3.6 which offers geometrical foundation for Ensemble machine learning. And we propose refinement and enhancement to both the theory and existing methods with the implementation of logifold structures for classification problems. The key new aspect of our implementation is the fuzzy domain of each model, which serves as a chart of the logifold. A trained model typically does not have a perfect accuracy rate and performs well only on certain part of the data, or for a subset of classes. A major step here is to find and record the domain of each model where it works well.
In classification problems, a dataset can be viewed as a fuzzy subset of $\mathbb{R}^{n}\times T$ for a finite set $T=\{c_{1},\ldots,c_{N}\}$ of classes. We restrict the domain of a model to $D\times T^{\prime}$ for some subsets $D\subset\mathbb{R}^{n}$ and $T^{\prime}\subset T$. Certainty scores are used to determine the (fuzzy) scopes of models, with the softmax function in the last layer of each node providing these scores. For each network model, we restrict the input domain to the subset of data that has certainty scores higher than a threshold. The threshold can be set based on performance on validation data. This produces a logifold which has a specific fuzzy domain, namely the union of the fuzzy domains of its charts.
2. Linear logical functions and their fuzzy analogs
Given a subset of $\mathbb{R}^{n}$, one would like to describe it as the zero locus, the image, or the graph of a function in a certain type. In analysis, we typically think of continuous/smooth/analytic functions. However, when the subset is not open, smoothness may not be the most relevant condition.
The success of network models has taught us a new kind of functions that are surprisingly powerful in describing datasets. Here, we formulate them using directed graphs and call them linear logical functions. First, they have the advantage of being logically interpretable in theory. Second, they are close analogues of quantum processes. Namely, they are made up of linear functions and certain non-linear activation functions, which are analogous to unitary evolution and quantum measurements. Third, it is natural to add fuzziness to these functions and hence they are better adapted to describe statistical data.
2.1. Linear logical functions and their graphs
We consider functions $D\to T$ for $D\subset\mathbb{R}^{n}$ and a finite set $T$ constructed from a graph as follows. Let $G$ be a finite directed graph that has no oriented cycle and has exactly one source vertex which has no incoming arrow and $|T|$ target vertices. Each vertex that has more than one outgoing arrows is equipped with an affine linear function $l=(l_{1},\ldots,l_{k})$ on $\mathbb{R}^{n}$, where the outgoing arrows at this vertex are one-to-one corresponding to the chambers in $\mathbb{R}^{n}$ subdivided by the hyperplanes $\{l_{i}=0\}$. (For theoretical purpose, we define these chambers to contain some of their boundary strata in a way such that they are disjoint and their union equals $\mathbb{R}^{n}$.)
Definition 2.1.
A linear logical function $f_{G,L}:D\to T$ is a function made in the following way from $(G,L)$, where $G$ is a finite directed graph that has no oriented cycle and has exactly one source vertex and $|T|$ target vertices,
[…]
$l_{v}$ are affine linear functions whose chambers in $D$ are one-to-one corresponding to the outgoing arrows of $v$. $(G,L)$ is called a linear logical graph.
Given $x\in D$, we get a path from the source vertex to one of the target vertices in $G$ as follows. We start with the source vertex. At a vertex, if there is only one outgoing arrow, we simply follow that arrow to reach the next vertex. If there are more than one outgoing arrow, we consider the chambers made by the affine linear function $l$ at that vertex, and pick the outgoing arrow that corresponds to the chamber that $x$ lies in. Since the graph is finite and has no oriented cycle, we will stop at a target vertex, which is associated to an element $t\in T$. This defines the function $f_{G,L}$ by setting $f_{G,L}(x)=t$.
Example 2.2.
Consider a feed-forward network model whose activation function at each hidden layer is the step function, and that at the last layer is the index-max function. The function is of the form
$\sigma\circ L_{N}\circ s_{N-1}\circ L_{N-1}\circ\ldots\circ s_{1}\circ L_{1}$
where $L_{i}:\mathbb{R}^{n_{i-1}}\to\mathbb{R}^{n_{i}}$ are affine linear functions with $n_{0}=n$, $s_{i}$ are the entrywise step functions and $\sigma$ is the index-max function. We make the generic assumption that the hyperplanes defined by $L_{2},\ldots,L_{N}$ do not contain the images of $s_{1},\ldots,s_{N-1}$. We claim that this is a linear logical function with target $T=\{1,\ldots,n_{N}\}$ (on any $D\subset\mathbb{R}^{n}$ where $\mathbb{R}^{n}$ is the domain of $L_{1}$).
The linear logical graph $(G,L)$ is constructed as follows. The source vertex $v_{0}$ is equipped with the affine linear function $L_{1}$. Then we make $N$ number of outgoing arrows of $v_{0}$ (and corresponding vertices) where $N$ is the number of chambers of $L_{1}$, which are one-to-one corresponding to the possible outcomes of $s_{1}$ (which form a finite subset of ${\{0,1\}}^{n_{1}}$). Then we consider $s_{2}\circ L_{2}$ restricted to this finite set, which also has a finite number of possible outcomes. This produces exactly one outgoing arrow for each of the vertices in the first layer. We proceed inductively. The last layer $\sigma\circ L_{N}$ is similar and has $n_{N}$ possible outcomes. Thus we obtain $(G,L)$ as claimed, where $L$ consists of only one affine linear function $L_{1}$ over the source vertex. See Figure 3.
Figure 3.
Example 2.3.
Consider a feed-forward network model whose activation function at each hidden layer is the ReLu function, and that at the last layer is the index-max function. The function takes the form
$\sigma\circ L_{N}\circ r_{N-1}\circ L_{N-1}\circ\ldots\circ r_{1}\circ L_{1}$
where $L_{i}:\mathbb{R}^{n_{i-1}}\to\mathbb{R}^{n_{i}}$ are affine linear functions with $n_{0}=n$, $r_{i}$ are the entrywise ReLu functions and $\sigma$ is the index-max function. We construct a linear logical graph $(G,L)$ which produces this function.
The first step is similar to the above example. Namely, the source vertex $v_{0}$ is equipped with the affine linear function $L_{1}$. Next, we make $N$ number of outgoing arrows of $v_{0}$ (and corresponding vertices), where $N$ is the number of chambers of $L_{1}$, which are one-to-one corresponding to the possible outcomes of the sign vector of $r_{1}$ (which form a finite subset of ${\{0,+\}}^{n_{1}}$). Now we consider the next linear function $L_{2}$. For each of these vertices in the first layer, we consider $L_{2}\circ r_{1}\circ L_{1}$ restricted to the corresponding chamber, which is a linear function on the original domain $R^{n}$, and we equip this function to the vertex. Again, we make a number of outgoing arrows that correspond to the chambers in $R^{n}$ made by this linear function. We proceed inductively, and get to the layer of vertices that correspond to the chambers of $L_{N-1}\circ r_{N-2}\circ L_{N-3}\circ\ldots\circ r_{1}\circ L_{1}$. Write $L_{N}=(l_{1},\ldots,l_{n_{N}})$, and consider $\tilde{L}_{N}=(l_{i}-l_{j}:i<j)$. At each of these vertices,
$\tilde{L}_{N}\circ r_{N-1}\circ L_{N-1}\circ\ldots\circ r_{1}\circ L_{1}$
restricted on the corresponding chamber is a linear function on the original domain $\mathbb{R}^{n}$, and we equip this function to the vertex and make outgoing arrows corresponding to the chambers of the function. In each chamber, the index $i$ that maximizes $l_{i}\circ r_{N-1}\circ L_{N-1}\circ\ldots\circ r_{1}\circ L_{1}$ is determined, and we make one outgoing arrow from the corresponding vertex to the target vertex $i\in T$. See Figure 4.
Figure 4.
In classification problems, $T$ is the set of labels for elements in $D$, and the data determines a subset in $D\times T$ as a graph of a function. Deep learning of network models provide a way to approximate the subset as the graph of a linear logical function $\mathrm{gr}(f_{G,L})$. Theoretically, this gives an interpretation of the set, namely, the linear logical graph gives a logical way to deduce the labels based on linear conditional statements on $D$.
The following lemma concerns about the monoidal structure on the set of linear logical functions on $D$.
Lemma 2.4.
Let $f_{G_{i},L_{i}}:D\to T_{i}$ be linear logical functions for $i\in I$ where $I=\{1,\ldots,k\}$. Then
$(f_{G_{i},L_{i}}:i\in I):D\to\prod_{i\in I}T_{i}$
is also a linear logical function.
Proof.
We construct a linear logical graph out of $G_{i}$ for $i\in I$ as follows. First, take the graph $G_{1}$. For each target vertex of $G_{1}$, we equip it with the linear function at the source vertex of $G_{2}$, and attach to it the graph $G_{2}$. The target vertices of the resulting graph are labeled by $T_{1}\times T_{2}$. Similarly, each target vertex of this graph is equipped with the linear function at the source vertex of $G_{3}$ and attached with the graph $G_{3}$. Inductively, we obtain the required graph, whose target vertices are labeled by $\prod_{i\in I}T_{i}$. By this construction, the corresponding function is $(f_{G_{i},L_{i}}:i\in I)$. ∎
$f_{(G,L)}$ admits the following algebraic expression in the form of a sum over paths which has an important interpretation in physics. The proof is straightforward and is omitted. A path in a directed graph is a finite sequence of composable arrows. The set of all linear combinations of paths and the trivial paths at vertices form an algebra by concatenation of paths.
Proposition 2.5.
Given a linear logical graph $(G,L)$,
$f_{(G,L)}(x)=h\left(\sum_{\gamma}c_{\gamma}(x)\,\gamma\right)=\sum_{\gamma}c_{\gamma}(x)\,h(\gamma)$
where the sum is over all possible paths $\gamma$ in $G$ from the source vertex to one of the target vertices; $h(\gamma)$ denotes the target vertex that $\gamma$ heads to; for $\gamma=a_{r}\ldots a_{1}$,
$c_{\gamma}(x)=\prod_{i=1}^{r}s_{a_{i}}(x)$
where $s_{a}(x)=1$ if $x$ lies in the chamber corresponding to the arrow $a$, or 00 otherwise. In the above sum, exactly one of the terms is non-zero.
2.2. Zero locus
Alternatively, we can formulate the graphs $\mathrm{gr}(f_{G,L})$ as zero loci of linear logical functions targeted at the field $\mathbb{F}_{2}$ with two elements as follows. Such a formulation has the advantage of making the framework of algebraic geometry available in this setting.
Proposition 2.6.
For each linear logical function $f_{G,L}:D\to T$, there exists a linear logical function $f_{\tilde{G},\tilde{L}}:D\times T\to\mathbb{F}_{2}$ whose zero locus in $D\times T$ equals $\mathrm{gr}(f_{G,L})$.
Proof.
Given a linear logical function $f_{G,L}:D\to T$, we construct another linear logical function $f_{\tilde{G},\tilde{L}}:D\times T\to\mathbb{F}_{2}$ as follows. Without loss of generality, let $T=\{1,\ldots,p\}$, so that $D\times T$ is embedded as a subset of $\mathbb{R}^{n+1}$. Any linear function on $\mathbb{R}^{n}$ is pulled back as a linear function on $\mathbb{R}^{n+1}$ by the standard projection $\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ that forgets the last component. Then $f_{G,L}$ is lifted as a linear logical function $D\times T\to T$.
Consider the corresponding graph $(G,L)$. For the $k$-th target vertex of $(G,L)$ (that corresponds to $k\in T$), we equip it with the linear function
$(y-(k-1/2),(k+1/2)-y):\mathbb{R}^{n+1}\to\mathbb{R}^{2}$
where $y$ is the last coordinate of $\mathbb{R}^{n+1}$. This linear function produces three chambers in $\mathbb{R}^{n+1}$. Correspondingly we make three outgoing arrows of the vertex. Finally, the outcome vertex that corresponds to (+,+)(+,+)( + , + ) is connected to the vertex $0\in\mathbb{F}_{2}$; the other two outcome vertices are connected to the vertex $1\in\mathbb{F}_{2}$. We obtain a linear logical graph $(\tilde{G},\tilde{L})$ and the corresponding function $f_{\tilde{G},\tilde{L}}:D\times T\to\mathbb{F}_{2}$.
By construction, $f_{\tilde{G},\tilde{L}}(x,y)=0$ for $(x,y)\in D\times T$ if and only if $y=f_{G,L}(x)\in T$. Thus, the zero locus of $f_{\tilde{G},\tilde{L}}$ is the graph of $f_{G,L}$. ∎
The set of functions (with a fixed domain) valued in $\mathbb{F}_{2}$ forms a unital commutative and associative algebra over $\mathbb{F}_{2}$, which is known as a Boolean algebra.
Proposition 2.7.
The subset of linear logical functions $D\to\mathbb{F}_{2}$ forms a Boolean ring $\mathcal{L}$ (for a fixed $D\subset\mathbb{R}^{n}$).
Proof.
We need to show that the subset is closed under addition and multiplication induced from the corresponding operations of $\mathbb{F}_{2}$.
Let $f_{(G_{1},L_{1})}$ and $f_{(G_{2},L_{2})}$ be linear logical functions $D\to\mathbb{F}_{2}$. By Lemma 2.4, $(f_{(G_{1},L_{1})},f_{(G_{2},L_{2})}):D\to{(\mathbb{F}_{2})}^{2}$ is a linear logical function. Consider the corresponding logical graph. The target vertices are labeled by $(s_{1},s_{2})\in{(\mathbb{F}_{2})}^{2}$. We connect each of them to the vertex $s_{1}+s_{2}\in\mathbb{F}_{2}$ by an arrow. This gives a linear logical graph whose corresponding function is $f_{(G_{1},L_{1})}+f_{(G_{2},L_{2})}$. We obtain $f_{(G_{1},L_{1})}\cdot f_{(G_{2},L_{2})}$ in a similar way. ∎
In this algebro-geometric formulation, the zero locus of $f_{G,L}:D\to\mathbb{F}_{2}$ corresponds to the ideal $(f_{G,L})\subset\mathcal{L}$.
2.3. Parametrization
Another geometric way to describe a set $X$ is via a parametrization $f:D\to X$. This way is commonly used in generative models in machine learning.
The graph $\mathrm{gr}(f_{G,L})$ of a linear logical function can also be put in parametric form. For the moment, let assume the domain $D$ to be finite. First, we need the following lemma.
Lemma 2.8.
Assume $D\subset\mathbb{R}^{n}$ is finite. Then the identity function $I_{D}:D\to D$ is a linear logical function.
Proof.
Since $D$ is finite, there exists a linear function $l:\mathbb{R}^{n}\to\mathbb{R}^{N}$ such that each chamber of $l$ contains at most one point of $D$. Then we construct the linear logical graph $G$ as follows. The source vertex is equipped with the linear function $l$ and outgoing arrows corresponding to the chambers of $l$. Elements of $D$ are identified as the target vertices of these arrows which correspond to chambers that contain them. The corresponding function $f_{G,L}$ equals $I_{D}$. ∎
Proposition 2.9.
Given a linear logical function $f_{G,L}:D\to T$ with finite $|D|$, there exists an injective linear logical function $D\to D\times T$ whose image equals $\mathrm{gr}(f_{G,L})$.
Proof.
By Lemma 2.8 and 2.4, $(I_{D},f_{G,L})$ is a linear logical function. By definition, its image equals $\mathrm{gr}(f_{G,L})$. ∎
2.4. Fuzzy linear logical functions
Another important feature of a dataset is its fuzziness. Below, we formulate the notion of a fuzzy linear logical function and consider its graph. Basic notions of fuzzy logic can be found in textbooks such as [18]. There are many developed applications of fuzzy logic such as modeling, control, pattern recognition and networks, see for instance [30, 25, 26, 28, 11].
Definition 2.10.
Let $G$ be a finite directed graph that has no oriented cycle, has exactly one source vertex and target vertices $t_{1},\ldots,t_{K}$ as in Definition 2.1. Each vertex $v$ of $G$ is equipped with a product of standard simplices
$P_{v}=\prod_{k=1}^{m_{v}}S^{d_{v,k}},\,\,\,S^{d_{v,k}}=\left\{(y_{0},\ldots,y_{d_{v,k}})\in\mathbb{R}_{\geq 0}^{d_{v,k}+1}:\sum_{i=0}^{d_{v,k}}y_{i}=1\right\}$
for some integers $m_{v}>0$, $d_{v,k}\geq 0$. $P_{v}$ is called the internal state space of the vertex $v$. Let $D$ be a subset of the internal state space of the source vertex of $G$. Each vertex $v$ that has more than one outgoing arrows is equipped with an affine linear function
$l_{v}:\prod_{k=1}^{m_{v}}\mathbb{R}^{d_{v,k}}\to\mathbb{R}^{j}$
for some $j>0$, whose chambers in the product simplex $P_{v}$ are one-to-one corresponding to the outgoing arrows of $v$. (In above, $\mathbb{R}^{d_{v,k}}$ is identified with the affine subspace $\left\{\sum_{i=0}^{d_{v,k}}y_{i}=1\right\}$ that contains $S^{d_{v,k}}$.) Let $L$ denote the collection of these linear functions. Moreover, each arrow $a$ is equipped with a continuous function
$p_{a}:P_{s(a)}\to P_{t(a)}$
where $s(a),t(a)$ denote the source and target vertices respectively.
We call $(G,L,P,p)$ a fuzzy linear logical graph. $(G,L,P,p)$ determines a function
$f_{(G,L,P,p)}:D\to P^{\textrm{out}}:=\coprod_{l=1}^{K}P_{t_{l}}$
as follows. Given $x\in D$, the collection $L$ of linear functions over vertices of $G$ evaluated at the image of $x$ under the arrow maps $p_{a}$ determines a path from the source vertex to one of the target vertices $t_{l}$. By composing the corresponding arrow maps $p_{a}$ on the internal state spaces along the path and evaluating at $x$, we obtain a value $f_{(G,L,P,p)}(x)\in P_{t_{l}}$. The resulting function $f_{(G,L,P,p)}$ is called a fuzzy linear logical function.
Remark 2.11.
Note that the linear functions $l_{v}$ in $L$ in the above definition have domain to be the internal state spaces over the corresponding vertices $v$. In comparison, the linear functions $l_{v}$ in $L$ in Definition 2.1 have domain to be the input space $\mathbb{R}^{n}$. A linear logical graph $(G,L)$ in Definition 2.1 has no internal state space except the $\mathbb{R}^{n}$ at the input vertex.
To relate the two notions given by the above definition and Definition 2.1, we can set $P_{v}=S^{n}$ to be the same for all vertices $v$ except the target vertices $t_{1},\ldots,t_{K}$, which are equipped with the zero-dimensional simplex (a point), and set $p_{a}$ to be identity maps for all arrows $a$ that are not targeted at any of $t_{i}$. Then $f_{(G,L,P,p)}$ reduces back to a linear logical function in Definition 2.1.
We call the corners of the convex set $P_{v}$ to be state vertices, which takes the form $e_{I}=(e_{i_{1}},\ldots,e_{i_{m_{v}}})\in P_{v}$ for a multi-index $I=(i_{1},\ldots,i_{m_{v}})$, where $\{e_{0},\ldots,e_{d_{v,k}}\}\subset\mathbb{R}^{d_{v,k}+1}$ is the standard basis. We can construct a bigger graph by replacing each vertex of $G$ by the collection of state vertices, and each arrow of $G$ by the collection of all possible arrows from source state vertices to target state vertices. Then the vertices of $G$ are interpreted as ‘layers’ or ‘clusters’ of vertices of this bigger graph. The input state $x\in D$, the arrow linear functions $L$ and the maps between state spaces $p$ determine the probability of getting to each target state vertex from the source vertex.
Under this interpretation, we take the target set to be the disjoint union of corners of $P_{t_{l}}$ at the target vertices $t_{1},\ldots,t_{K}$:
(2.1) $T=\coprod_{l=1}^{K}T_{l}:=\coprod_{l=1}^{K}\left\{e_{I}:I=(i_{1},\ldots,i_{m_{t_{l}}})\textrm{ for }i_{k}\in\{0,\ldots,d_{t_{l},k}\}\right\}$
which is a finite set. The function $f=f_{(G,L,P,p)}$ determines the probability of the outcome for each input state $x\in D$ as follows. Let $f(x)\in P_{t_{l}}=\prod_{k=1}^{m_{t_{l}}}S^{d_{t_{l},k}}$ for some $l=1,\ldots,K$. Then the probability of being in $T_{j}$ is zero for $j\not=l$. Writing $f(x)=(f_{1}(x),\ldots,f_{m_{t_{l}}}(x))$ for $f_{k}(x)\in S^{d_{t_{l},k}}$, the probability of the output to be $t=e_{I}\in T_{l}$ for $I=(i_{1},\ldots,i_{m_{t_{l}}})$ is given by $\prod_{k=1}^{m_{t_{l}}}f^{(i_{k})}_{k}(x)\in[0,1]$.
Example 2.12.
Consider a feed-forward network model whose activation function at each hidden layer is the sigmoid function $\tilde{s}$, and that at the last layer is the softmax function $\tilde{\sigma}$:
$f=\tilde{\sigma}\circ L_{N}\circ\tilde{s}_{N-1}\circ L_{N-1}\circ\ldots\circ\tilde{s}_{1}\circ L_{1}.$
We set $(G,L,P,p)$ as follows. $G$ is the graph that has $(N+1)$ vertices $v_{0},\ldots,v_{N}$ with arrows $a_{i}$ from $v_{i-1}$ to $v_{i}$ for $i=1,\ldots,N$. $L$ is just an empty set. $P_{i}:={(S^{1})}^{m_{i}}$ for $i=0,\ldots,N-1$, where $m_{i}$ is the dimension of the domain of $L_{i+1}$. The one-dimensional simplex $S^{1}$ is identified with the interval $[0,1]$. $P_{N}:=S^{m_{N}}$ where $m_{N}$ is the dimension of the target of $L_{N}$. Then
$p_{i}:=\tilde{s}_{i}\circ L_{i}|_{{[0,1]}^{m_{i-1}}}\textrm{ for}i=1,\ldots,N-1$
$p_{N}:=\tilde{\sigma}\circ L_{N}|_{{[0,1]}^{m_{N-1}}}.$
Then $f=f_{(G,L,P,p)}$.
Example 2.13.
Consider a feed-forward network model whose activation function at each hidden layer is the ReLu function, and that at the last layer is the softmax function $\tilde{\sigma}$:
$f=\tilde{\sigma}\circ L_{N}\circ r_{N-1}\circ L_{N-1}\circ\ldots\circ r_{1}\circ L_{1}.$
We take $G$ to be the logical graph constructed in Example 2.3 (Figure 4) with the last two layers of vertices replaced by a single target vertex $t$. We take the internal state space to be $P_{v}:={(S^{1})}^{n}$ at every vertex $v$ except the target vertex, whose internal state space is defined to be the simplex $P_{t}:=S^{d}$ where $d$ is the target dimension of $L_{N}$. $p_{a}$ is defined to be the identity function for all arrows except those that target at $t$. $L$ is the taken to be the collection of linear functions on $\mathbb{R}^{n}$ over the vertices as in Example 2.3, except the vertices that are adjacent to the last target vertex $t$ and $t$ itself, which do not have more than one outgoing arrows. On each adjacent vertex of the target vertex, recall that $L_{N}\circ r_{N-1}\circ L_{N-1}\circ\ldots\circ r_{1}\circ L_{1}$ restricted on the corresponding chamber is a linear function $l$ on $\mathbb{R}^{n}$. Then $p_{a}$ for the corresponding arrow $a$ is defined to be $\tilde{\sigma}\circ l$. By this construction, we have $f=f_{(G,L,P,p)}$.
As in Proposition 2.5, $f_{(G,L,P,p)}$ can be expressed in the form of sum over paths.
Proposition 2.14.
Given a fuzzy linear logical graph $(G,L,P,p)$,
$f_{(G,L,P,p)}(x)=\sum_{\gamma}c_{\gamma}(x)\,p_{\gamma}(x)$
where the sum is over all possible paths $\gamma$ in $G$ from the source vertex to one of the target vertices; for $\gamma=a_{r}\ldots a_{1}$, $p_{\gamma}(x)=\prod_{i=1}^{r}p_{a_{r}}\ldots p_{a_{1}}(x)$;
$c_{\gamma}(x)=\prod_{i=1}^{r}s_{a_{i}}(p_{a_{i-1}\ldots a_{1}}(x))$
where $s_{a}(x)=1$ if $x\in P_{t_{a}}$ lies in the chamber corresponding to the arrow $a$, or 00 otherwise. In the above sum, exactly one of the terms is non-zero.
2.5. As a fuzzy subset
A fuzzy subset of a topological measure space $X$ is a continuous measurable function $\mathcal{F}:X\to[0,1]$. This generalizes the characteristic function of a subset. The interval $[0,1]$ can be equipped with the semiring structure whose addition and multiplication are taking maximum and minimum respectively. This induces a semiring structure on the collection of fuzzy subsets $\mathcal{F}:X\to[0,1]$ that plays the role of union and intersection operations.
The graph $\mathrm{gr}(f)$ of a function
$f:D\to\coprod_{l=1}^{K}P_{t_{l}},$
where $P_{t_{l}}$ are products of simplices as in Definition 2.10, is defined to be the fuzzy subset in $D\times T$, where $T$ is defined by Equation (2.1), given by the probability at every $(x,t)\in D\times T$ determined by $f$ in Remark 2.11.
The following is a fuzzy analog of Proposition 2.6.
Proposition 2.15.
Let $f_{(G,L,P,p)}$ be a fuzzy linear logical function. Let $\mathcal{F}:D\times T\to[0,1]$ be the characteristic function of its graph where $T$ is defined by (2.1). Then $\mathcal{F}$ is also a fuzzy linear logical function.
Proof.
Similar to the proof of Proposition 2.6, we embed the finite set $T$ as the subset $\{1,\ldots,|T|\}\subset\mathbb{R}$. The affine linear functions on $\mathbb{R}^{n}\supset D$ in the collection $L$ are pulled back as affine linear functions on $\mathbb{R}^{n+1}\supset D\times T$. Similarly, for the input vertex $v_{0}$, we replace the product simplex $P_{v_{0}}$ by $\tilde{P}_{v_{0}}:=P_{v_{0}}\times[0,|T|+1]$ (where the interval $[0,|T|+1]$ is identified with $S^{1}$); for the arrows $a$ tailing at $v_{0}$, $p_{a}$ are pulled back to be functions $\tilde{P}_{v_{0}}\to P_{h(a)}$. Then we obtain $(\tilde{L},\tilde{P},\tilde{p})$ on $G$.
For each of the target vertices $t_{l}$ of $G$, we equip it with the linear function
$(y-3/2,y-5/2,\ldots,y-(|T|-1/2)):\mathbb{R}^{n+1}\to\mathbb{R}^{|T|-1}$
where $y$ is the last coordinate of the domain $\mathbb{R}^{n+1}$. It divides $\mathbb{R}^{n+1}$ into $|T|$ chambers that contain $\mathbb{R}^{n}\times\{j\}$ for some $j=1,\ldots,|T|$. Correspondingly we make $|T|$ outgoing arrows of the vertex $t_{l}$. The new vertices are equipped with the internal state space $P_{t_{l}}$, and the new arrows are equipped with the identity function $P_{t_{l}}\to P_{t_{l}}$. Then we get $|T|K$ additional vertices, where $K$ is the number of target vertices $t_{l}$ of $G$. Let’s label these vertices by $v_{j,l}$ for $j\in T$ and $l\in\{1,\ldots,K\}$. Each of these vertices are connected to the new output vertex by a new arrow. The new output vertex is equipped with the internal state space $S^{1}\cong[0,1]$. The arrow from $v_{j,l}$ to the output vertex is equipped with the following function $p_{j,l}:P_{v_{j,l}}\to[0,1]$. If $j\not\in T_{l}$, then we set $p_{j,l}\equiv 0$. Otherwise, for $j=e_{I}\in T_{l}$ and $I=(i_{1},\ldots,i_{m_{t_{l}}})$, $p_{j,l}:=\prod_{k=1}^{m_{t_{l}}}u^{(i_{k})}_{k}$ where $u_{k}^{(0)},\ldots,u_{k}^{(d_{t_{l},k})}$ are the coordinates of $S^{d_{t_{l},k}}\subset\mathbb{R}_{\geq 0}^{d_{t_{l},k}+1}$. This gives the fuzzy linear logical graph $(\tilde{G},\tilde{L},\tilde{P},\tilde{p})$ whose associated function $f_{(\tilde{G},\tilde{L},\tilde{P},\tilde{p})}:D\times T\to[0,1]$ is the characteristic function. ∎
The above motivates us to consider fuzzy subsets whose characteristic functions are fuzzy linear logical functions $\mathcal{F}:X\to[0,1]$. Below, we will show that they form a sub-semiring, that is, they are closed under fuzzy union and intersection. We will need the following lemma analogous to Lemma 2.4.
Lemma 2.16.
Let $f_{G_{i},L_{i},P_{i},p_{i}}:D\to P^{\textrm{out}}_{i}$ be fuzzy linear logical functions for $i\in I$ where $I=\{1,\ldots,k\}$, and assume that the input state space $P_{i,\mathrm{in}}$ are the same for all $i$. Then
$(f_{G_{i},L_{i},P_{i},p_{i}}:i\in I):D\to\prod_{i\in I}P^{\textrm{out}}_{i}$
is also a fuzzy linear logical function.
Proof.
By the proof of Lemma 2.4, we obtain a new graph $(G,L)$ from $(G_{i},L_{i})$ for $i=1,\ldots,k$ by attaching $(G_{i+1},L_{i+1})$ to the target vertices of $(G_{i},L_{i})$. For the internal state spaces, we change as follows. First, we make a new input vertex $\tilde{v}_{0}$ and an arrow $\tilde{a}_{0}$ from $\tilde{v}_{0}$ to the original input vertex $v_{0}$ of $(G,L)$. We denote the resulting graph by $(\tilde{G},\tilde{L})$. We define $\tilde{P}_{\tilde{v}_{0}}:=P_{v_{0}}$, $\tilde{P}_{v_{0}}:=\prod_{i=1}^{k}P_{v_{0}}$ where $P_{v_{0}}=P_{i,\mathrm{in}}$ for all $i$ by assumption, and $\tilde{p}_{\tilde{a}_{0}}:P_{v_{0}}\to\prod_{i=1}^{k}P_{v_{0}}$ to be the diagonal map $\tilde{p}_{\tilde{a}_{0}}=(\mathrm{Id},\ldots,\mathrm{Id})$. The internal state spaces $P_{v}$ over vertices $v$ of $G_{1}$ are replaced by $P_{v}\times\prod_{i=2}^{k}P_{v_{0}}$, and $p_{a}$ for arrows $a$ of $G_{1}$ are replaced by $(p_{a},\mathrm{Id},\ldots,\mathrm{Id})$. Next, over vertices $v$ of the graph $G_{2}$ that is attached to the target vertex $t_{l}$ of $G_{1}$, the internal state space $P_{v}$ is replaced by $P_{t_{l}}\times P_{v}\times\prod_{i=3}^{k}P_{v_{0}}$, and $p_{a}$ for arrows $a$ of $G_{2}$ are replaced by $(\mathrm{Id},p_{a},\mathrm{Id},\ldots,\mathrm{Id})$. Inductively, we obtain the desired graph $(\tilde{G},\tilde{L},\tilde{P},\tilde{p})$. ∎
Proposition 2.17.
Suppose $\mathcal{F}_{1},\mathcal{F}_{2}:X\to[0,1]$ are fuzzy subsets defined by fuzzy linear logical functions. Then $\mathcal{F}_{1}\cup\mathcal{F}_{2}$ and $\mathcal{F}_{1}\cap\mathcal{F}_{2}$ are also fuzzy subsets defined by fuzzy linear logical functions.
Proof.
By the previous lemma, $(\mathcal{F}_{1},\mathcal{F}_{2})=f_{(G,L,P,p)}:X\to[0,1]\times[0,1]$ for some fuzzy linear logical graph $(G,L,P,p)$, which has a single output vertex whose internal state space is $[0,1]\times[0,1]\cong S^{1}\times S^{1}$. We attach an arrow $a$ to this output vertex. Over the new target vertex $v$, $P_{v}:=[0,1]$; $p_{a}:=\max:[0,1]\times[0,1]\to[0,1]$ (or $p_{a}:=\min$). Then we obtain $(\tilde{G},\tilde{L},\tilde{P},\tilde{p})$ whose corresponding fuzzy function defines $\mathcal{F}_{1}\cup\mathcal{F}_{2}$ (or $\mathcal{F}_{1}\cap\mathcal{F}_{2}$ respectively). ∎
Remark 2.18.
For $f:P^{\textrm{in}}\to P^{\textrm{out}}$ where $P^{\textrm{in}},P^{\textrm{out}}$ are product simplices, we can have various interpretations.
  1. (1)
    As a usual function, its graph is in the product $P^{\textrm{in}}\times P^{\textrm{out}}$.
  2. (2)
    As a fuzzy function on $P^{\textrm{in}}$: $P^{\textrm{in}}\to T$ where $T$ is the finite set of vertices of the product simplex $P^{\textrm{out}}$, its graph is a fuzzy subset in $P^{\textrm{in}}\times T$.
  3. (3)
    The domain product simplex $P^{\textrm{in}}$ can also be understood as a collection of fuzzy points over $V$, the finite set of vertices of $P^{\textrm{in}}$, where a fuzzy point here just refers to a probability distribution (which integrates to $1$).
  4. (4)
    Similarly, $P^{\textrm{in}}\times P^{\textrm{out}}$ can be understood as a collection of fuzzy points over $V\times T$. Thus, the (usual) graph of $f$ can be interpreted as a sub-collection of fuzzy points over $V\times T$.
$(\mathrm{Id},f)$ gives a parametric description of the graph of a function $f$. The following ensures that it is a fuzzy linear logical function if $f$ is.
Corollary 2.19.
Let $f:D\to P^{\mathrm{out}}$ be a fuzzy linear logical function. Then $(\mathrm{Id},f):D\to D\times P^{\mathrm{out}}$ is also a fuzzy linear logical function whose image is the graph of $f$.
Proof.
By Lemma 2.16, it suffices to know that $\mathrm{Id}:D\to D$ is a fuzzy linear logical function. This is obvious: we take the graph with two vertices serving as input and output, which are connected by one arrow. The input and output vertices are equipped with the internal state spaces that contain $D$, and $p$ is just defined by the identity function. ∎
Remark 2.20.
Generative deep learning models widely used nowadays can be understood as parametric descriptions of data sets $X$ by fuzzy linear logical functions $f:D\to X$ (where $D$ and $X$ are embedded in certain product simplices $P^{\mathrm{in}}$ and $P^{\mathrm{out}}$ respectively).
2.6. Non-linearity in a quantum-classical system
The fuzzy linear logical functions in Definition 2.10 have the following quantum analog. Quantum systems and quantum random walks are well known and studied, see for instance [5] and [24]. On the other hand, they depend only linearly on the initial state in probability. The motivation of this subsection is to investigate how non-linear dependence on the initial probability distribution can come up.
Definition 2.21.
Let $G$ be a finite directed graph that has no oriented cycle and has exactly one source vertex and target vertices $t_{1},\ldots,t_{K}$. Each vertex $v$ is equipped with a product of projectifications of Hilbert spaces over complex numbers:
$Q_{v}:=\prod_{l=1}^{m_{v}}\mathbb{P}(H_{v,l})$
for some integer $m_{v}>0$. We fix an orthonormal basis in each Hilbert space $H_{v,l}$, which gives a basis in the tensor product:
$E^{(v)}=\left\{e^{(v)}_{I}=(e^{(v)}_{i_{1}},\ldots,e^{(v)}_{i_{m_{v}}}):e^{(v)}_{i_{l}}\textrm{ is a basic vector of }H_{v,l}\right\}.$
For each vertex $v$ that has more than one outgoing arrows, we make a choice of a decomposition of the set $E^{(v)}$ into subsets that are one-to-one corresponding to the outgoing arrows. Each arrow $a$ is equipped with a map $q_{a}$ from the corresponding subset of basic vectors $e^{(t(a))}_{I}$ to $Q_{h(a)}$.
Let’s call the tuple $(G,Q,E,q)$ to be a quantum logical graph.
We obtain a probabilistic map $f_{(G,Q,E,q)}:Q^{\mathrm{in}}\to T$ as follows. Given a state $\vec{w}=(w_{1},\ldots,w_{m_{v}})\in Q_{v}$ at a vertex $v$, we make a quantum measurement and $\vec{w}$ projects to one of the basic elements $e^{(v)}_{I}$ with probability $\prod_{l=1}^{m_{v}}|\langle w_{l},e^{(v)}_{i_{l}}\rangle|^{2}$. The outcome $e^{(v)}_{I}$ determines which outgoing arrow $a$ to pick, and the corresponding map $q_{a}$ sends it to an element of $Q_{h(a)}$. Inductively we obtain an element $f_{(G,Q,E,q)}(w)\in T$.
However, such a process is simply linearly depending on the initial condition in probability: the probabilities of outcomes of the quantum process $f_{(G,Q,E,q)}(w)$ for an input state $w$ (which is complex-valued) simply linearly depends on the modulus of components of the input $w$. In other words, the output probabilities are simply obtained by a matrix multiplication on the input probabilities. To produce non-linear physical phenomena, we need the following extra ingredient.
Let’s consider the state space $\mathbb{P}^{n}$ of a single particle. A basis gives a map $\mu:\mathbb{P}^{n}\to S^{n}$ to the simplex $S^{n}$ (also known as the moment map of a corresponding torus action on $\mathbb{P}^{n}$):
$\mu=\left(\frac{|z_{0}|^{2}}{|z_{0}|^{2}+\ldots+|z_{n}|^{2}},\ldots,\frac{|z_{n}|^{2}}{|z_{0}|^{2}+\ldots+|z_{n}|^{2}}\right):\mathbb{P}^{n}\to S^{n}.$
The components of the moment map are the probability of quantum projection to basic states of the particle upon observation. By the law of large numbers, if we make independent observations of particles in an identical quantum state $\vec{z}=[z_{0}:\ldots:z_{n}]\in\mathbb{P}^{n}$ for $N$ times, the average of the observed results (which are elements in $\{(1,0,\ldots,0),\ldots,(0,\ldots,0,1)\}$) converges to $\mu(\vec{z})\in S^{n}$ as $N\to\infty$.
The additional ingredient we need is a choice of a map $s:S^{n}\to\mathbb{P}^{m}$ and $m,n\in\mathbb{Z}_{>0}$. For instance, for $m=n=1$, we set the initial phase of the electron spin state according to a number in $[0,1]$.
Upon an observation of a state, we obtain a point in $\{(1,0,\ldots,0),\ldots,(0,\ldots,0,1)\}\subset S^{n}$. Now if we have $N$ particles simultaneously observed, we obtain $N$ values, whose average is again a point $p$ in the simplex $S^{n}$. By $s$, these are turned to $N$ quantum particles in state $s(p)\in\mathbb{P}^{m}$ again.
$\mu:\mathbb{P}^{n}\to S^{n}$ and $s:S^{n}\to\mathbb{P}^{m}$ give an interplay between quantum processes and classical processes with averaging. Averaging in the classical world is the main ingredient to produce non-linearity from the linear quantum process.
Now, let’s modify Definition 2.21 by using $s:S^{n}\to\mathbb{P}^{m}$. Let $P_{v}$ be the product simplex corresponding to $Q_{v}$ at each vertex. Moreover, as in Definition 2.1 and 2.10 for (fuzzy) linear logical functions, we equip each vertex with affine linear functions $l_{v}$ whose corresponding systems of inequalities divide $P_{v}$ into chambers. This decomposition of $P_{v}$ plays the role of the decomposition of $E^{(v)}$ in Definition 2.21. The outgoing arrows at $v$ are in a one-to-one correspondence with the chambers. Each outgoing arrow $a$ at $v$ is equipped with a map $\tilde{q}_{a}$ from the corresponding chamber of $P_{v}$ to $Q_{h(a)}$. $\tilde{q}_{a}$ can be understood as an extension of $q_{a}$ (whose domain is a subset of corners of $P_{t(a)}$) in Definition 2.21.
Definition 2.22.
We call the tuple $(G,Q,E,L,\tilde{q})$ (where $L$ is the collection of affine linear functions $l_{v}$) to be a quantum-classical logical graph.
Given $N$ copies of the same state $\vec{w}$ in $Q_{v}$, we first take a quantum projection of these and they become elements in $P_{v}$. We take an average of these $N$ elements, which lies in a certain chamber defined by $l_{v}$. The chamber corresponds to an outgoing arrow $a$, and the map $\tilde{q}_{a}$ produces $N$ elements in $Q_{h(a)}$. Inductively we obtain a quantum-classical process $Q^{\mathrm{in}}\to T$.
For $N=1$, this essentially produces the same linear probabilistic outcomes as in Definition 2.21. On the other hand, when $N>1$, the process is no longer linear and produces a fuzzy linear logical function $P^{\mathrm{in}}\to P^{\mathrm{out}}$.
In summary, non-linear dependence on the initial state results from averaging of observed states.
Remark 2.23.
We can allow loops or cycles in the above definition. Then the system may run without stop. In this situation, the main object of concern is the resulting (possibly infinite) sequence of pairs $(v,s)$, where $v$ is a vertex of $G$ and $s$ is a state in $Q_{v}$. This gives a quantum-classical walk on the graph $G$.
We can make a similar generalization for (fuzzy) linear logical functions by allowing loops or cycles. This is typical in applications in time-dependent network models.
3. Linear logical structures for a measure space
In the previous section, we have defined linear logical functions based on a directed graph. In this section, we will first show the equivalence between our definition of linear logical functions and semilinear functions[31] in the literature. Thus, the linear logical graph we have defined can be understood as a representation of semilinear functions. Moreover, fuzzy and quantum logical functions that we define can be understood as deformations of semilinear functions.
Next, we will consider measurable functions and show that they can be approximated and covered by semilinear functions. This motivates the definition of a logifold, which is a measure space that has graphs of linear logical functions as local models.
3.1. Equivalence with semilinear functions
Let’s first recall the definition of semilinear sets.
Definition 3.1.
For any positive integer $n$, semilinear sets are the subsets of $\mathbb{R}^{n}$ that are finite unions of sets of the form
(3.1) $\{x\in\mathbb{R}^{n}:f_{1}(x)=\cdots=f_{k}(x)=0,g_{1}(x)>0,\ldots,g_{l}(x)>0\}$
where the $f_{i}$ and $g_{j}$ are affine linear functions.
A function $f:D\to T$ on $D\subset\mathbb{R}^{n}$, where $T$ is a discrete set, is called to be semilinear if for every $t\in T$, $f^{-1}\{t\}$ equals to the intersection of $D$ with a semilinear set.
Now let’s consider linear logical functions defined in the last section. We show that the two notions are equivalent (when the target set is finite). Thus, a linear logical graph can be understood as a graphical representation (which is not unique) of a semi-linear function. From this perspective, the last section provides fuzzy and quantum deformations of semi-linear functions.
Theorem 3.2.
Consider $f:D\to T$ for a finite set $T=\{t_{1},\ldots,t_{s}\}$ where $D\subset\mathbb{R}^{n}$. $f$ is a semilinear function if and only if it is a linear logical function.
Proof.
It suffices to consider the case $D=\mathbb{R}^{n}$. We will use the following terminologies for convenience. Let $V(G)$ and $E(G)$ be the sets of vertices and arrows respectively for a directed graph $G$. A vertex $v\in V(G)$ is called to be nontrivial if it has more than one outgoing arrows. It is said to be simple if it has exactly one outgoing arrow. We call a vertex that has no outgoing arrow to be a target, and that has no incoming arrow to be a source. For a target $t$, let $R_{t}$ be the set of all paths from the source to target $t$.
Consider a linear logical function $f=f_{(G,L)}:\mathbb{R}^{n}\to T$. Let $p$ be a path in $R_{t}$ for some $t\in T$. Let $\{v_{1},\ldots,v_{k}\}$ be the set of non-trivial vertices that $p$ passes through. This is a non-empty set unless $f$ is just a constant function (recall that $G$ has only one source vertex). At each of these vertices $v_{i}$, $\mathbb{R}^{n}$ is subdivided according to the affine linear functions $g_{i,1},\ldots,g_{i,N_{i}}$ into chambers $C_{i,1},\ldots,C_{i,m_{i}}$ where $m_{i}$ is the number of its outgoing arrows. All the chambers $C_{i,j}$ are semilinear sets.
For each path $p\in R_{t}$, we define a set $E_{p}$ such that $x\in E_{p}$ if $x$ follows path $p$ to get the target $t$. Then $E_{p}$ can be represented as $C_{1,j_{1}}\cap\cdots\cap C_{k,j_{k}},$ which is semilinear. Moreover, the finite union
$f^{-1}(t)=\bigcup_{p\in R_{t}}E_{p}$
is also a semilinear set. This shows that $f$ is a semilinear function.
Conversely, suppose that we are given a semilinear function. Without loss of generality, we can assume that $f$ is surjective. For every $t\in T$, $f^{-1}(t_{i})$ is a semilinear set defined by a collection of affine linear functions in the form of (3.1). Let =$\mathcal{F}=\{l_{1},\ldots,l_{N}\}$ be the union of these collections over all $t\in T$.
Now we construct a linear logical graph associated to $f$. We consider the chambers made by $(l_{1},-l_{1},\ldots,l_{N},-l_{N})$ by taking the intersection of the half spaces $l_{i}\geq 0$, $l_{i}<0$, $-l_{i}\geq 0$, $-l_{i}<0$. We construct outgoing arrows of the source vertex associated with these chambers.
For each $t\in T$, $l_{j}$ occurs in defining $f^{-1}(t)$ as either one of the following ways:
  1. (1)
    $l_{j}>0$, which is equivalent to $l_{j}\geq 0$ and $-l_{j}<0$,
  2. (2)
    $l_{j}=0$, which is equivalent to $l_{j}\geq 0$ and $-l_{j}\geq 0$
  3. (3)
    $l_{j}$ is not involved in defining $f^{-1}(t)$.
Thus, $f^{-1}(t)$ is a union of a sub-collection of chambers associated to the outgoing arrows. Then we assign these outgoing arrows with the target vertex $t$. This is well-defined since $f^{-1}(t)$ for different $t$ are disjoint to each other. Moreover, since $\bigcup_{t}f^{-1}(t)=\mathbb{R}^{n}$, every outgoing arrow is associated with a certain target vertex.
In summary, we have constructed a linear logical graph $G$ which produces the function $f$.
The above equivalence between semilinear functions and linear logical functions naturally generalizes to definable functions of other types of $o$-minimal structures. They provide the simplest class of examples in $o$-minimal structures for semialgebraic and subanalytic geometry [31]. Topology of sub-level sets of definable functions was recently investigated in [22]. Let’s first recall the basic definitions.
Definition 3.3.
[31] A structure $\mathcal{S}$ on $\mathbb{R}$ consists of a Boolean algebra $\mathcal{S}_{n}$ of subsets of $\mathbb{R}^{n}$ for each $n=0,1,2,\ldots,$ such that
  1. (1)
    the diagonals $\{x\in\mathbb{R}^{n}:x_{i}=x_{j}\},1\leq i<j\leq n$ belong to $\mathcal{S}_{n}$;
  2. (2)
    $A\in\mathcal{S}_{m},B\in\mathcal{S}_{n}\implies A\times B\in\mathcal{S}_{m+n}$;
  3. (3)
    $A\in\mathcal{S}_{n+1}\implies\pi(A)\in\mathcal{S}_{n}$, where $\pi:\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ is the projection map defined by $\pi(x_{1},\ldots,x_{n+1})=(x_{1},\ldots,x_{n})$;
  4. (4)
    the ordering $\{(x,y)\in\mathbb{R}^{2}:x<y\}$ of $\mathbb{R}$ belongs to $\mathcal{S}_{2}$.
A structure $\mathcal{S}$ is $o$-minimal if the sets in $\mathcal{S}_{1}$ are exactly the subsets of $\mathbb{R}$ that have only finitely many connected components, that is, the finite unions of intervals and points.
Given a collection $\mathcal{A}$ of subsets of the Cartesian spaces $\mathbb{R}^{n}$ for various $n$, such that the ordering $\{(x,y):x<y\}$ belongs to $\mathcal{A}$, define $\operatorname{Def}(\mathcal{A})$ as the smallest structure on the real line containing $\mathcal{A}$ by adding the diagonals to $\mathcal{A}$ and closing off under Boolean operations, cartesian products and projections. Sets in $\operatorname{Def}(\mathcal{A})$ are said to be definable from $\mathcal{A}$ or just definable if $\mathcal{A}$ is clear from context.
Given definable sets $A\subset\mathbb{R}^{m}$ and $B\subset\mathbb{R}^{n}$ we say that a map $f:A\to B$ is definable if its graph $\Gamma(f)=\{(x,f(x))\in\mathbb{R}^{m+n}:x\in A\}$ is definable.
Remark 3.4.
If $\mathcal{A}$ consist of the ordering, the singletons $\{r\}$ for any $r\in\mathbb{R}$, the graph in $\mathbb{R}^{2}$ of scalar multiplications $x\mapsto\lambda x:\mathbb{R}\to\mathbb{R}$ for any $\lambda\in\mathbb{R}$, and the graph of addition $\{(x,y,z)\in\mathbb{R}^{3}:z=x+y\}$. Then $\operatorname{Def}(\mathcal{A})$ consists of semilinear sets for various positive integer $n$ (Definition 3.1).
Similarly, if $\mathcal{A}$ consist of the ordering, singletons, and the graphs of addition and multiplication, then $\operatorname{Def}(\mathcal{A})$ consists of semi-algebraic sets, which are finite unions of sets of the form
$\left\{x\in\mathbb{R}^{n}:f(x)=0,g_{1}(x)>0,\ldots,g_{l}(x)>0\right\}$
where $f$ and $g_{1},\ldots,g_{l}$ are real polynomials in $n$ variables, due to the Tarski-Seidenberg Theorem[6].
One obtains semi-analytic sets in which the above $f,g_{1},\ldots,g_{l}$ become real analytic functions by including graphs of analytic functions. Let an be the collection $\mathcal{A}$ and of the functions $f:\mathbb{R}^{n}\to\mathbb{R}$ for all positive integer $n$ such that $f|_{I^{n}}$ is analytic, $I=[-1,1]\subset\mathbb{R}$, and $f$ is identically 00 outside the cubes. The theory of semi-analytic sets and subanalytic sets show that $\operatorname{Def}(\textbf{an})$ is $o$-minimal, and relatively compact semi-analytic sets have only finitely many connected components. See [6] for efficient exposition of the Łojasiewicz-Gabrielov-Hironaka theory of semi- and subanalytic sets.
Theorem 3.5.
Let’s replace the collection of affine linear functions at vertices in Definition 2.1 by polynomials and call the resulting functions to be polynomial logical functions. Then $f:D\to T$ for a finite set $T$ is a semi-algebraic function if and only if $f$ is polynomial logical functions.
The proof of the above theorem is similar to that of Theorem 3.2 and hence omitted.
3.2. Approximation of measurable functions by linear logical functions
We consider measurable functions $f:D\to T$, where $0<\mu\left(D\right)<\infty$ and $T$ is a finite set. The following approximation theorem for measurable functions has two distinct features since $T$ is a finite set. First, the functions under consideration, and linear logical functions that we use, are discontinuous. Second, the ‘approximating function’ actually exactly equals to the target function in a large part of $D$. Compared to traditional approximation methods, linear logical functions have an advantage of being representable by logical graphs, which have fuzzy or quantum generalizations.
Theorem 3.6 (Universal approximation theorem for measurable functions).
Let $\mu$ be the standard Lebesgue measure on $\mathbb{R}^{n}$. Let $f:D\to T$ be a measurable function with $\mu(D)<\infty$ and a finite target set $T$. For any $\epsilon>0$, there exists a linear logical function $L:D\to\widetilde{T}$, where $\widetilde{T}=T\cup\{\ast\}$ is $T$ adjunct with a singleton, and a measurable set $E\subset D$ with $\mu(E)<\epsilon$ such that $L|_{D-E}\equiv f|_{D-E}$.
Proof.
Let :=$\mathcal{R}:=\left\{\prod_{k=1}^{n}(a_{k},b_{k}]\subset\mathbb{R}^{n}:a_{k}<b_{k}\textrm{ for all k}\right\}$ be the family of rectangles in $\mathbb{R}^{n}$. We will use the well-known fact that for any measurable set $U$ of finite Lebesgue measure, there exists a finite subcollection $\left\{R_{j}:j=1,\ldots,N\right\}$ of $\mathcal{R}$ such that $\mu\left(U\,\triangle\,\bigcup_{1}^{N}R_{j}\right)<\epsilon$ (see for instance [15]). Here, $A\,\triangle\,B:=(A-B)\cup(B-A)$ denotes the symmetric difference of two subsets $A,B$.
Suppose that a measurable function $f:D\to T=\{t_{1},\ldots,t_{m}\}$ and $\epsilon>0$ be given. For each $t\in T$, let $\mathcal{S}_{t}$ be a union of finitely many rectangles of $\mathcal{R}$ that approximates $f^{-1}(t)\subset D$ (that has finite measure) in the sense that $\mu\left((f^{-1}(t)\,\triangle\,\mathcal{S}_{t})\right)<\frac{\epsilon}{|T|}$. Note that $\mathcal{S}_{t}$ is a semilinear set.
The case $m=1$ is trivial. Suppose $m>1$. Define semilinear sets $\mathcal{S}_{\ast}=\underset{i<j}{\bigcup}\left(\mathcal{S}_{t_{i}}\cap\mathcal{S}_{t_{j}}\right)$ and $\mathfrak{S}_{t}=\mathcal{S}_{t}\setminus\mathcal{S}_{\ast}$ for each $t\in T$. Now we define $L:D\to\widetilde{T}$,
$L(p)=\begin{cases}t&\text{if }p\in\mathfrak{S}_{t}\cap D\\ \ast&\text{if }p\in D\setminus\underset{t\in T}{\bigcup}\mathfrak{S}_{t}\end{cases}$
which is a semilinear function on $D$.
If $p\in\mathcal{S}_{\ast}\cap D$ then $p\in\mathcal{S}_{t_{i}}\cap\mathcal{S}_{t_{j}}$ for some $t_{i},t_{j}\in T$ with $t_{i}\neq t_{j}$. In such a case, $p\in f^{-1}(t_{i})$ implies $p\notin f^{-1}(t_{j})$. It shows $\mathcal{S}_{\ast}\cap D\subset\underset{t\in T}{\bigcup}\left(\mathcal{S}_{t}\setminus f^{-1}(t)\right)$. Also we have $D\setminus\underset{t\in T}{\bigcup}\mathcal{S}_{t}\subset\underset{t\in T}{\bigcup}\left(f^{-1}(t)\setminus\mathcal{S}_{t}\right)$. Therefore
$D\setminus\underset{t\in T}{\bigcup}\mathfrak{S}_{t}=\left(D\setminus\underset{t\in T}{\bigcup}\mathcal{S}_{t}\right)\cup\left(D\cap\mathcal{S}_{\ast}\right)\subset\underset{t\in T}{\bigcup}f^{-1}(t)\,\triangle\,\mathcal{S}_{t}$
and hence
$\mu\left(D\setminus\underset{t\in T}{\bigcup}\mathfrak{S}_{t}\right)\leq\sum_{t\in T}\mu\left(f^{-1}(t)\,\triangle\,\mathcal{S}_{t}\right)<\epsilon.$
By Theorem 3.2, $L$ is a linear logical function. ∎
Corollary 3.7.
[…]
3.3. Linear logifold
To be more flexible, we can work with Hausdorff measure which is recalled as follows.
Definition 3.8.
Let $p\geq 0$, $\delta>0$. For any $U\subset\mathbb{R}^{n}$, $\operatorname{Diam}(U)$ denotes the diameter of $U$ defined by the supremum of distance of any two points in $U$. For a subset $E\subset\mathbb{R}^{n}$, define
$H^{p}_{\delta}\left(E\right):=\underset{\mathcal{U}_{\delta}(E)}{\inf}\,\sum_{U\in\mathcal{U}_{\delta}(E)}\operatorname{Diam}\left(U\right)^{p}$
[…]
Definition 3.9.
A linear logifold is a pair $(X,\mathcal{U})$, where $X$ is a topological space equipped with a $\sigma$-algebra and a measure $\mu$, $\mathcal{U}$ is a collection of pairs $(U_{i},\phi_{i})$ where $U_{i}$ are subsets of $X$ such that $\mu(U_{i})>0$ and $\mu(X-\bigcup_{i}U_{i})=0$; $\phi_{i}$ are measure-preserving homeomorphisms between $U_{i}$ and the graphs of linear logical functions $f_{i}:D_{i}\to T_{i}$ (with an induced Hausdorff measure), where $D_{i}\subset\mathbb{R}^{n_{i}}$ are $H^{p_{i}}$-measurable subsets in certain dimension $p_{i}$, and $T_{i}$ are discrete sets.
The elements of $\mathcal{U}$ are called charts. A chart $(U,\phi)$ is called to be entire up to measure $\epsilon$ if $\mu(X-U)<\epsilon$.
Comparing to a topological manifold, we require $\mu(U_{i})>0$ in place of openness condition. Local models are now taken to be graphs of linear logical functions in place of open subsets of Euclidean spaces.
Then the results in the last subsection can be rephrased as follows.
Corollary 3.10.
Let $f:D\to T$ be a measurable function on a measurable set $D\subset\mathbb{R}^{n}$ of finite measure with a finite target set $T$. For any $\epsilon>0$, its graph $\mathrm{gr}(f)\subset D\times T$ can be equipped with a linear logifold structure that has an entire chart up to measure $\epsilon$.
Remark 3.11.
In[4], relations between neural networks and quiver representations were studied. In [20, 21], a network model is formulated as a framed quiver representation; learning of the model was formulated as a stochastic gradient descent over the corresponding moduli space. In this language, we now take several quivers, and we glue their representations together (in a non-linear way) to form a ‘logifold’.
Remark 3.12.
Topological change under different scales is the beautiful subject of persistent homology [14, 33, 8]. Moreover, there are also interesting categorical or sheaf theoretical approaches (see for instance [10, 16]) to understand relations between different components of a system.
In a similar manner, we define a fuzzy linear logifold below. By Remark 2.11 and (2) of Remark 2.18, a fuzzy linear logical function has a graph as a fuzzy subset of $D\times T$. We are going to use the fuzzy graph as a local model for a fuzzy space $(X,\mathcal{P})$.
Definition 3.13.
A fuzzy linear logifold is a tuple $(X,\mathcal{P},\mathcal{U})$, where
  1. (1)
    $X$ is a topological space equipped with a measure $\mu$;
  2. (2)
    $\mathcal{P}:X\to[0,1]$ is a continuous measurable function;
  3. (3)
    $\mathcal{U}$ is a collection of tuples $(\rho_{i},\phi_{i},f_{i})$, where $\rho_{i}$ are measurable functions $\rho_{i}:X\to[0,1]$ with $\sum_{i}\rho_{i}\leq 1_{X}$ that describe fuzzy subsets of $X$, whose supports are denoted by by $U_{i}=\{x\in X:\rho_{i}(x)>0\}\subset X$;
    $\phi_{i}:U_{i}\to D_{i}\times T_{i}$
    are measure-preserving homeomorphisms where $T_{i}$ are finite sets in the form of (2.1) and $D_{i}\subset\mathbb{R}^{n_{i}}$ are $H^{p_{i}}$-measurable subsets in certain dimension $p_{i}$; $f_{i}$ are fuzzy linear logical functions on $D_{i}$ whose target sets are $T_{i}$ described in Remark 2.11;
  4. (4)
    the induced fuzzy graphs $\mathcal{F}_{i}:D_{i}\times T_{i}\to[0,1]$ of $f_{i}$ satisfy
    (3.2) $\mathcal{P}=\sum_{i}\rho_{i}\cdot\phi_{i}^{*}(\mathcal{F}_{i}).$
In applications to classification problems, we set $X=\mathbb{R}^{n}\times T$ for a finite set $T$, and $\mathcal{P}:X\to[0,1]$ describe how likely an element of $\mathbb{R}^{n}\times T$ is classified as ‘yes’. In this situation $\mathcal{P}$ should satisfy $\sum_{t\in T}\mathcal{P}(s,t)\leq 1$ for every $s\in\mathbb{R}^{n}$. $U_{i}$ can simply be taken to be $D_{i}\times T_{i}$ where $D_{i}\subset\mathbb{R}^{n}$ is measurable and $T_{i}\subset T$; $\phi_{i}$ is just taken to be identity. $\rho_{i}$ can be understood as the weight (or confidence) of the corresponding linear logical interpretation.
Remark 3.14.
We allow $T_{i}$ to be a proper subset of $T$ and the collection of $U_{i}$ is not required to cover the whole $X$. The fuzzy set $(X,\mathcal{P})$ has been covered as long as Equation (3.2) is satisfied.
Remark 3.15.
The fuzzy linear logical functions $\phi_{i}$ and $\phi_{j}$ for different $i,j$ are not required to agree on the common intersections $D_{i}\cap D_{j}$. In applications, it often happens that two individuals come up with different answers in classifying the same input. The equation 3.2 provides a method for combining multiple ‘models’ in Ensemble theory.
If we have a notion of morphism between logifolds, we can have isomorphism classes among logifolds, and try to classify them using persistent homology through filterations indexed by fuzziness. We will develope this approach in future.
4. Ensemble learning and logifolds
In this section, we briefly review ensemble learning [27] and make a rigorous mathematical formulation ofensemble methods via logifolds. Moreover, we define fuzzy domains and introduce a more refined voting method which can improve the accuracy of predictions made by the ensemble.
4.1. A brief review of Ensemble theory
Ensemble machine learning (multiple classifer system, or ensemble systmes) utilize more than one number of classifiers to make decisions. Dasarathy and Sheela[12] were early contributors to this theory, proposing the partitioning feature space using multiple classifiers. There are several reasons to use ensemble system, such as smoothing decision boundary, reducing bias of classifiers, or addressing issues related to data volume. A widely accepted key to a successful ensemble system is achieving diversity among classifiers. [9, 13, 32] provide good reviews of this theory.
Broadly speaking, designing an ensemble system involves determining how to obtain classifiers with diversity and how to combine their predictions effectively. Here, we briefly introduce popular methods Bagging[7], short for bootstrap aggregating trains multiple classifiers, each on a randomly sampled subset of the training dataset. Boosting, such as AdaBoost(Adaptive Boosting)[29, 17] itertatively trains classifiers by focusing on the instances they misclassified in previous rounds. In Mixture of Experts[19], each classifier specializes in different tasks or subsets of dataset, with a ‘gating’ layer, which determines weights for the combination of classifiers.
Given multiple classifiers, an ensemble system makes decisions based on predictions from diverse classifiers, which may produce different answers for the same instance. Therefore, a rule for combining predictions is necessary. If each classifier provides a discrete answer, voting strategies , such as Borda count or simple majority voting, can be applied. If a measure of generalization performance is available for each classifier, or if classifiers output continuous values that represent certainty or confidence, a weighted sum can be applied based on these measures. Alternatively, the combining weights can be trained. Kuncheva introduced a decision profile matrix to combine continuous-valued predictions, with methods such as the mean rule, min/max rule, or weighted average[1]. Different combination rules may yield different outcomes from the same ensemble system for the same instance.
In this section, we introduce our combination rules for multiple classifiers, which uses fuzzy domains and models with different target classes. We view each trained model as a fuzzy linear logical function, forming a chart of fuzzy linear logifold structure. The domain of each model may be a proper subset of its feature space and the entire set of target classes. For each trained model, a fuzzy domain is defined, with its fuzziness measured by the certainty for each instance. The model provides an answer when it is likely to have the correct classification for that instance and the scope of classes it is proficient at classifying.
4.2. Mathematical Description of Neural Network Learning
Let $X\subset\mathbb{R}^{n}$ and $T=\{t_{1},\ldots,t_{N}\}$, and $\mathcal{Z}\subset X\times T$ is given, which we take as a data sample. The space $X$ is typically referred to as the feature space, while each $t_{i}\in T$ represents a class. We embed $T$ as the corners of the standard simplex $S^{N-1}\subset\mathbb{R}^{N}$, and hence $\mathcal{Z}$ is embedded in $\mathbb{R}^{n}\times\mathbb{R}^{N}$. One wants to use functions $X\to S^{N-1}$ produced by neural networks to interpolate the subset $\mathcal{Z}$. A deep neural network and its function are mathematically defined as follows.
Definition 4.1.
The underlying graph of a neural network is a finite directed graph $G$. Each vertex $v$ is associated with $\mathbb{R}^{n_{v}}$ for some $n_{v}\in\mathbb{Z}_{>0}$, together with a non-linear function $\mathbb{R}^{n_{v}}\to\mathbb{R}^{n_{v}}$ called an activation function.
Let $\Theta$ be the vector space of linear representations. A linear representation associates each arrow $a$ with a linear map $\mathbb{R}^{n_{s(a)}}\to\mathbb{R}^{n_{t(a)}}$, where $s(a),t(a)$ are the source and target vertices respectively.
Let’s fix $\gamma$ to be a linear combination of paths between two fixed vertices $s$ and $t$ in $G$. The associated network function $f_{\theta}:\mathbb{R}^{n_{s}}\to\mathbb{R}^{n_{t}}$ for each $\theta\in\Theta$ is defined to be the corresponding function obtained by the sum of compositions of linear functions and activation functions along the paths of $\gamma$.
One would like to minimize the function $C_{\mathcal{Z}}:\Theta\to\mathbb{R}$,
$C_{\mathcal{Z}}(\theta):=\sum_{(x,y)\in\mathcal{Z}}\|f_{\theta}(x)-y\|^{2}$
which measures the distance between the graph of $f_{\theta}$ and $\mathcal{Z}$. To do this, one takes a stochastic gradient descent(SGD) over $\Theta$. In a discrete setting, it is given by the following equation:
$\theta_{k+1}=\theta_{k}-\eta\nabla C_{\mathcal{Z}}-\eta W_{k}$
where $\eta\in R_{>0}$ is called step size or learning rate, $W_{k}$ is the $k^{\textrm{th}}$ noise or Brownian Motion, and $\nabla C_{\mathcal{Z}}$ denotes the gradient vector field of $C_{\mathcal{Z}}$. (In practice, the sample $\mathcal{Z}$ is divided into batches and $C$ is a sum over a batch.)
For practical purpose, the completion of the computational process is marked by the verification of epochs. Then the hyper-parameter space $\mathcal{H}$ for SGD is a subspace $\{(\eta,\textrm{Batch size},\textrm{Epochs},\textrm{Noise})\}\subset\mathbb{R}^{3}\times\mathcal{D}$, where $\mathcal{D}$ is the space of $\mathbb{R}$-valued random variables with zero mean and finite variance. This process is called the training procedure, and the function obtained is called a trained model.
Remark 4.2.
Bagging[7] is a method used to produce trained models by selecting different subsets of the training dataset $\mathcal{Z}$ during training procedure. Boosting[29], on the other hand, iteratively trains model on a dataset $\mathcal{Z}^{\prime}$, where the cost function, such as $C_{\mathcal{Z}}$, has not been well-optimized.
4.2.1. Logifold structure of an ensemble
A trained model $g$ is a fuzzy linear logical function, denoted $g_{(G,L,P,p)}$, where $G$ is the neural network with one target vertex, $L$ and $p$ are the affine maps and activation functions, and $g:X\to S^{N-1}$ maps $x$ to $\left(g_{1}(x),\ldots,g_{N}(x)\right)$. This implies that each classifier in an ensemble system can be interpreted as a tuple within the collection $\mathcal{U}$, as defined in Definition 3.13.
Let $g_{i}:X\to S^{N-1}$ be trained models, and let $G_{i}:X\times T\to[0,1]$ be the graphs induced by $g_{i}$ for $i=1,\ldots,k$. The predictions made by $\{g_{1},\ldots,g_{k}\}$ are combined via a weighted sum with weights $\rho_{i}:X\to[0,1]$. Then we have $P:X\times T\to[0,1]$ such that
$P(x,t)=\sum_{i=1}^{k}\rho_{i}(x)G_{i}(x,t),$
and a collection of tuples $\mathcal{U}=\left(\rho_{i},\textrm{Id}_{X\times T},g_{i}\right)$ equipped with the induced fuzziness $G_{i}:X\times T\to[0,1]$. The triple $\left(X,P,\mathcal{U}\right)$ forms a fuzzy linear logifold.
Remark 4.3.
The weight $\rho_{i}$ can be chosen based on the generalization performance of the trained model, such as the accuracy evaluated on a validation dataset. Each $\rho_{i}$ corresponding to the $M_{i}$ determines the rule for combining predictions made by $M_{i}$.
In this setup, trained models now serve as charts of our logifold. The main new ingredient is that, for each model, we define a fuzzy domain and restrict the model to this domain using a ‘certainty score’. In other words, for a given data point, each model is required to ‘vote’ only when it is certain about the prediction. Additionally, we allow modelsto focus on partial target sets. Thus different models may operate over distinct domains and targets, which is a special feature of our logifold formulation.
4.2.2. Specialization
A trained model $f$ can be turned into a ‘specialist’ for a subset of target classes $T_{i}=\{t_{1},\ldots,t_{m}\}\subsetneq T=\{c_{1},\ldots,c_{N}\}$. Let $G$ be the underlying graph of a neural network, and let $\gamma$ denote the linear combinations of paths $\gamma_{1},\ldots,\gamma_{k}$ from $s$ to $t$ in $G$ such that the associated network function $f_{\theta}$ is the trained model $f$.
[…]
In following sections, we introduce the methods and implementation detail for ‘finding the fuzzy domain’ and ‘the refined voting system’.
4.3. Notions in practice
We assume that our problem is a classification task, with classes denoted as $T=\{c_{1},\ldots,c_{N}\}$. Let $X\subset\mathbb{R}^{n}$ be a measurable space, and let $S^{N-1}$ be the $(N-1)$-dimensional standard simplex in $\mathbb{R}^{N}$. For any measurable subset $A$ of $X$, we denote the measure of $A$ by $\left|A\right|$.
Definition 4.4.
Let $T$ be a finite set and $\widetilde{T}$ be a subset of $\mathcal{P}\left(T\right)$ where $\mathcal{P}\left(T\right)$ is the power set of $T$ We call $\widetilde{T}$ a set of logical targets if it is non-empty subset of $\mathcal{P}\left(T\right)\setminus\{\emptyset\}$, such that $t\cap t^{\prime}=\emptyset$ for all distinct $t,t^{\prime}\in\widetilde{T}$. An element of $\widetilde{T}$ is called a logical target.
  1. (1)
    A logical target $t\in\widetilde{T}$ is called thin (or fine) if it is a singleton and thick otherwise.
  2. (2)
    The union $\bigcup\widetilde{T}$ is called the flattening of $\widetilde{T}$, denoted by $\overline{T}$.
  3. (3)
    We say that $\widetilde{T}$ is full in $T$ if $T=\overline{T}$, and fine if it has no thick logical target.
Let $\mathcal{M}$ represent a system comprising various trained models $M_{i}$, modeled by neural networks and indexed by $i\in\mathcal{I}$. Each model $M_{i}$ has flexibility to select target classes for training, such that each target $t$ of $M_{i}$ is a non-empty subset of $T$. For instance, we can train a model on a training sample of the form $(x,t)$ where $t\subset T$ and $|t|\neq 1$.
Let $S_{i}$ denote the training sample for the training procedure of model $M_{i}$, and let $T_{i}$ denote the set of possible targets based on the training sample $S_{i}$. Specifically, we define
$T_{i}=\{t\subset T:(x,t)\in S_{i}\}.$
Thus, $T_{i}=\{t_{i,1},\ldots,t_{i,N_{i}}\}$ is the set of targets for model $M_{i}$, where $t_{i,j}\subset T$ and $t_{i,l}\cap t_{i,m}=\emptyset$ for all $t_{i,l}\neq t_{i,m}$. Therefore, $T_{i}$ is a set of logical targets, and the trained model $M_{i}$ is a function
$M_{i}:\mathcal{D}_{i}\to S^{N_{i}-1},$
where $\mathcal{D}_{i}$ is the domain of all possible inputs for $M_{i}$. We refer to $T_{i}$ as the set of logical target classes of the trained model $M_{i}$. A trained model $M_{i}$ is said to be full, or fine if $T_{i}$ is full, or fine, respectively.
Definition 4.5.
Let $g:X\to S^{n}$ be a function such that $g(x)={\left(g_{1}(x),\ldots,g_{n}(x)\right)}$ with positive integer $n$.
  1. (1)
    The function $G:X\times T\to[0,1]$ induced by $g$, mapping $(x,t_{j})\mapsto g_{j}(x)$, is called fuzziness of the graph of $g$.
  2. (2)
    $\arg\max g(x)$ and $\max g(x)$ are called the prediction and certainty of $g$ at $x$, respectively.
For each instance $x\in\mathcal{D}_{i}$, the output $M_{i}(x)\in S^{N_{i}-1}$ is referred to as the answer (or decision vector) produced by $M_{i}$. If $M_{i}(x)=(p_{i,1}(x),\ldots,p_{i,N_{i}}(x))$, then $\sum_{j=1}^{k}p_{i,j}(x)=1$, where $p_{i,j}(x)\in[0,1]$. Here, $p_{i,j}(x)$ can be interpreted as the certainty of $M_{i}$ regarding whether instance $x$ belongs to the logical target $t_{j}\in T_{i}$. Let $P_{i}(x)$ and $C_{i}(x)$ denote the prediction and certainty of $M_{i}$ at $x$, respectively.
Let us denote a dataset by $\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$ with input $\mathcal{X}\subset\mathcal{D}_{i}$ and output $\mathcal{Y}\subset T_{i}$. Subscripts can be added if necessary, such as $\mathcal{Z}_{\textrm{training}}$ or $\mathcal{Z}_{\textrm{validation}}$.
Given a dataset $\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$ and $\alpha\in[0,1]$, $Z_{\alpha,i}$ denotes the certain part with certainty threshold $\alpha$ of the model $M_{i}$ for $\mathcal{Z}$, which is defined as
$\mathcal{Z}_{\alpha,i}$ $=\left(\mathcal{X}_{\alpha,i},\mathcal{Y}_{\alpha,i}\right),$
$\mathcal{X}_{\alpha,i}$ $=\mathcal{X}\cap\{x\in\mathcal{X}:C_{i}(x)\geq\alpha\},$
$\mathcal{Y}_{\alpha,i}$ =$=\{y\in\mathcal{Y}:(x,y)\in\mathcal{Z}\textrm{ and }x\in\mathcal{X}_{\alpha,i}\}.$
The union $\mathcal{Z}_{\alpha}=\bigcup_{i\in\mathcal{I}}{\mathcal{Z}_{\alpha,i}}=\left(\mathcal{X}_{\alpha},\mathcal{Y}_{\alpha}\right)$ is called the certain part with certainty threshold $\alpha$ of $\mathcal{M}$ for $\mathcal{Z}$ Table 1 summarizes the notations introduced so far.
$\mathcal{M}$ A system of trained models $\{M_{i}\}_{i\in\mathcal{I}}$
$T=\{c_{1},\ldots,c_{N}\}$ The set of classes
$\mathcal{I}$ The set of indices of trained models in $\mathcal{M}$
$M_{i}$ A trained model with index $i$
$T_{i}=\{t_{1},\ldots,t_{N_{i}}\}$ and $\overline{T}_{i}=\bigcup T_{i}$ Set of logical targets of $M_{i}$ and its flattening
$M_{i}(x)$ The answer (or decision vector) of $M_{i}$ for $x$
$p_{i,j}(x)$ The certainty of $M_{i}$ whether $x\in t_{j}\in T_{i}$
$C_{i}(x)=\max M_{i}(x)$ The certainty of $M_{i}$ at $x$
$P_{i}(x)=\arg\max M(x)$ The prediction of $M_{i}$ at $x$
$\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$ A dataset
$\mathcal{Z}_{\alpha,i}=\left(\mathcal{X}_{\alpha,i},\mathcal{Y}_{\alpha,i}\right)$ The certain part of $\mathcal{Z}$ at certainty threshold $\alpha$ by $M_{i}$
$\mathcal{Z}_{\alpha}=\left(\mathcal{X}_{\alpha},\mathcal{Y}_{\alpha}\right)$ The certain part of $\mathcal{Z}$ at certainty threshold $\alpha$ by $\mathcal{M}$
Table 1. Table of frequently used symbols
4.4. Finding Fuzzy domains
Suppose that $M_{i}:\mathcal{D}_{i}\to T_{i}=\{t_{1},\ldots,t_{N_{i}}\}$ and a dataset $\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$ are given. We define the accuracy, coverage, and accuracy for each logical target $t_{j}\left(j=1,\ldots,N_{i}\right)$ of the trained model $M_{i}$. Each metric is a function of the certainty thresholds $\alpha\in[0,1]$ and is defined as follows:
  • The feature space of the certain part of $\mathcal{Z}$ at certainty threshold $\alpha$ by $M_{i}$:
    $\mathcal{X}_{i}(\alpha):=\mathcal{X}_{\alpha,i},$
    and therefore coverage of $M_{i}$ at $\alpha$ is
    $\textrm{Coverage}=\frac{\left|\mathcal{X}_{i}(\alpha)\right|}{\left|\mathcal{X}\right|}.$
    If the trained model is given by $M$, then $\mathcal{X}_{M}(\alpha)$ denotes the feature space of certain part of $\mathcal{Z}$ at certainty threshold $\alpha$ by $M$.
  • The accuracy of trained model $M_{i}$ over $\mathcal{X}_{i}(\alpha)$:
    $\Phi_{\mathcal{X},i}(\alpha):=\frac{\left|\{x\in\mathcal{X}_{\alpha,i}:\left(x,P_{i}(x)\right)\in\mathcal{Z}_{\alpha,i}\}\right|}{\left|\mathcal{X}_{\alpha,i}\right|}.$
  • The accuracy of trained model $M_{i}$ over $\mathcal{X}_{i}(\alpha)$ for a logical target $t_{j}\in T_{i}$:
    $\phi_{\mathcal{X},i,t_{j}}(\alpha):=\frac{\left|\{x\in\mathcal{X}_{\alpha,i}:\left(x,P_{i}(x)\right)\in\mathcal{Z}\textrm{ and }P_{i}(x)=t_{j}\}\right|}{\left|\{x\in\mathcal{X}_{\alpha,i}:P_{i}(x)=t_{j}\right|}.$
The fuzzy domain of $M_{i}$ is defined as the triple of functions:
$\left(\Phi_{\mathcal{X},i},\mathcal{X}_{i},(\phi_{\mathcal{X},i,t_{j}}:t_{j}\in T_{i})\right).$
When the index is not necessary, we suppress index of $M_{i}$ in each function for simplicity. In practice, we calculate and record the fuzzy domain for the validation dataset of each model $M$ in the system $\mathcal{M}$. Algorithm 1 in Appendix A summarizes the procedure for obtaining the fuzzy domain of a given trained model.
4.5. Voting system
We introduce a voting system that can incorporate trained models that are neither full nor fine, involving fuzzy domains. To do this, we fisrt define weighted answer according to the fuzzy domain of each trained model.
Let $\{M_{i}\}_{i\in\mathcal{I}}$ be a collection of trained models comprising $\mathcal{M}$ with their logical target classes $\{T_{i}\}_{i\in\mathcal{I}}$ and flattenings $\{\overline{T}_{i}\}_{i\in\mathcal{I}}$. Let $T=\bigcup_{i\in\mathcal{I}}\overline{T}_{i}$, denoted by $\{c_{1},\ldots,c_{N}\}$. For each $c\in T$ and each $M_{i}$, assuming each component in a non-fine target has the same weight, we can convert each logical target $t\in T_{i}$ into an $N$ dimensional vector as follows:
Definition 4.6.
Let $T=\{c_{1},\ldots,c_{N}\}$ be a finite set and $\widetilde{T}\subset\mathcal{P}(T)$ be a set of logical targets. For each $t\in\widetilde{T}$, define
$\vec{t}=(a_{1},\ldots,a_{N})\quad\textrm{where }a_{i}=\begin{cases}|t|^{-1}&\textrm{ if }c_{i}\in t,\\ 0&\textrm{ otherwise.}\end{cases}$
We call $\vec{t}$ the vectorized logical target for each $t\in T$.
Therefore, we can form an $N_{i}\times N$ matrix $\widehat{T}_{i}$ from $T_{i}$ such that
$\widehat{T}_{i}=\begin{bmatrix}\vec{t}_{i,1}\\ \vdots\\ \vec{t}_{i,N_{i}}\end{bmatrix}$
where $T_{i}=\{t_{i,1},\ldots,t_{i,N_{i}}\}$.
Given the fuzzy domain of $M_{i}$ on $\mathcal{X}$, each model $M_{i}$ has a triple of functions $\left(\Phi_{\mathcal{X},i},\mathcal{X}_{i},\left(\phi_{\mathcal{X},i,t}\right)_{t\in T_{i}}\right)$. Define a matrix-valued function $\mathfrak{P}_{i}:[0,1]\to\textrm{Mat}_{\left|\mathcal{X}\right|\times N_{i}}\left(\mathbb{R}\right)$ of certainty thresholds $\alpha$ such that
$\left[\mathfrak{P}_{i}(\alpha)\right]_{u,v}=\begin{cases*}\phi_{\mathcal{X},i,t_{i,v}}(\alpha)\quad\text{ if }P_{i}(x_{u})=t_{i,v}\text{ and }x_{u}\in\mathcal{X}_{\alpha,i}\\ \phi_{\mathcal{X},i,t_{i,v}}(0)\quad\text{ if }P_{i}(x_{u})=t_{i,v}\text{ and }x_{u}\in\mathcal{X}\setminus\mathcal{X}_{\alpha,i}\\ 0\quad\quad\quad\quad\quad\text{otherwise.}\end{cases*}$
The weighted answer of $M_{i}$ for $x_{u}\in\mathcal{X}$ to a class $c\in T$ at certainty threshold $\alpha$ on $\mathcal{X}$ is the $(u,c)-$ component of product of matrices
$\mathfrak{P}_{i}(\alpha)\cdot\widehat{T}_{i}.$
In brief, the matrix $\mathfrak{P}_{i}(\alpha)\cdot\widehat{T}_{i}$ is called the weighted answer of $M_{i}$ on $\mathcal{X}$ at certainty threshoold $\alpha$.
Example 4.7.
Let $M_{0}$ be a trained model within a system $\mathcal{M}$ with $T_{0}=\{\{c_{1}\},\{c_{2},c_{3}\}\}$ and $T=\{c_{1},c_{2},c_{3},c_{4}\}$. Suppose that given $\mathcal{Z}=\{(x_{1},y_{1}),(x_{2},y_{2})\}$ and $\mathcal{X}_{0}(0.5)=\{x_{1}\}$. The accuracies for each logical target at thresholds $\{0,0.5\}$ are given as:
$\left(\phi_{\mathcal{X},0,\{0\}}(0),\phi_{\mathcal{X},0,\{1,2\}}(0)\right)$ $=(0.65,0.6),$
$\left(\phi_{\mathcal{X},0,\{0\}}(0.5),\phi_{\mathcal{X},0,\{1,2\}}(0.5)\right)$ $=(0.9,0.7),$
by the fuzzy domain of $M_{0}$. Given that $P_{0}(x_{1})=\{0\}$ and $P_{0}(x_{2})=\{1,2\}$, we have
$\mathfrak{P}_{0}(0.5)=\begin{bmatrix}0.9&0\\ 0&0.6\end{bmatrix},\quad\widehat{T}_{i}=\begin{bmatrix}1&0&0&0\\ 0&\frac{1}{2}&\frac{1}{2}&0\end{bmatrix},$
and therefore, the weighted answer of $M_{0}$ at certainty threshold $0.5$ on $\mathcal{X}$ is
$\mathfrak{P}_{0}(0.5)\cdot\widehat{T}_{0}=\begin{bmatrix}0.9&0&0&0\\ 0&0.3&0.3&0\end{bmatrix}.$
4.5.1. Refined vote at a node of the target tree.
For a flattening of logical target $\overline{T}_{i}$, we can define a subset of indices $\mathcal{J}\subset\mathcal{I}$ as $\mathcal{J}=\{k\in\mathcal{I}:\overline{T}_{k}=\overline{T}_{i}\}$. We define the weighted answer for $x_{u}\in\mathcal{X}$ of trained models with the same flattening logical targets $\overline{T}_{i}$ as
$\sum_{i\in\mathcal{J}}\left(u^{\textrm{th}}\textrm{ row vector of }\mathfrak{P}_{i}(\alpha)\cdot\widehat{T}_{i}\right).$
This constitutes our voting system by ‘weighing accuracy’. Algorithm 2 in Appendix A provides the pseudo-algorithm for computing this weighted answer.
We assign a partial order to the system $\mathcal{M}$, where answers from each trained model accumulate according to this order. The partial order is encoded by a tree $\mathfrak{T}$, which we call a target tree.
If all trained models are full in $T$, then the target tree is trivial, consisting of a single node $s$. Let us assume that $\mathcal{M}$ includes both full and not-full trained models in $T$. Define $\mathcal{T}$ as the set of flattenings of logical targets, that is
$\mathcal{T}:=\left\{\overline{T}_{i}:i\in\mathcal{I}\right\},$
with $T\in\mathcal{T}$.
Place a root node and associate with $T$. Then construct a tree graph satisfying that for any flattenins of logical targets $\overline{T}_{i}\in\mathcal{T}$ and a node $v$ in the target tree associated with a flattening of targets $\mathcal{T}_{v}$, if $\overline{T}_{i}\subsetneq\mathcal{T}_{v}$, and there is no other $\overline{T}_{j}\in\mathcal{T}$ such that $\overline{T}_{i}\subsetneq\overline{T}_{j}\subsetneq\mathcal{T}_{v}$, then place node $u$ associated with $\overline{T}_{i}$ as a child of the parent node $v$.
Let $V$ denote the set of nodes in $\mathfrak{T}$. We define two functions $V\to\mathcal{P}(T)$ mapping $s\mapsto\mathcal{T}_{s}$, and $s\mapsto\mathcal{I}_{s}:=\{i\in\mathcal{I}:\overline{T}_{i}=\mathcal{T}_{s}\}$ recording the indices of the trained models corresponding to node $s$.
At a node $s$ in target tree $\mathfrak{T}$, the answer $\mathcal{A}$ for $x_{u}$ with certainty threshold $\alpha$ is defined as:
(4.1) $\mathcal{A}(s,\alpha,x_{u})=\sum_{i\in\mathcal{I}_{s}}\left(u^{\textrm{th}}\textrm{ row vector of }\mathfrak{P}_{i}(\alpha)\cdot\widehat{T}_{i}\right).$
The prediction for $x_{u}$ at node $s$ by $\mathcal{M}$ with certainty threshold $\alpha$ is given by $\arg\max\mathcal{A}(s,\alpha,x_{u})$.
4.5.2. Accumulation of votes along a path in the target tree.
If $\mathfrak{T}$ is trivial, the refined vote at the root according to the Equation 4.1 provides the answer. Let us assume that $\mathfrak{T}$ is nontrivial.
We define $\widetilde{\mathcal{A}}(s,\alpha,x)=\left(\widetilde{\mathcal{A}}(s,\alpha,x)_{1},\ldots,\widetilde{\mathcal{A}}(s,\alpha,x)_{N}\right)\in\mathbb{R}^{N}$ as the modified refined vote for $x$ by $\mathcal{M}$ at certainty threshold $\alpha$ and node $s$ in the target tree $\mathfrak{T}$ as follows:
(4.2) $\widetilde{\mathcal{A}}(s,\alpha,x)_{j}=\begin{cases}\mathcal{A}(s,\alpha,x)_{j}&\textrm{if }c_{j}\in\mathcal{T}_{s},\\ 0&\textrm{otherwise,}\end{cases}$
where $\mathcal{A}(s,\alpha,x_{u})=\left(\mathcal{A}(s,\alpha,x_{u})_{1},\ldots,\mathcal{A}(s,\alpha,x_{u})_{N}\right)\in\mathbb{R}^{N}.$ Note that $\widetilde{\mathcal{A}}(s,\alpha,x)=\mathcal{A}(s,\alpha,x)$ for the root node $s$.
For a given path $\gamma=(s_{1},s_{2},\ldots,s_{k})$ in $\mathfrak{T}$ such that $s_{1}$ is the root and $s_{i+1}$ is a child node of $s_{i}$ for $i=1,\ldots,k-1$, we define $\mathcal{X}(\gamma,\alpha)$ as the certain part of $\mathcal{M}$ along with the path $\gamma=(s_{1},\ldots,s_{k})$ in the target tree $\mathfrak{T}$ at certainty threshold $\alpha$ as
$\mathcal{X}(\gamma,\alpha):=\bigcup_{i=1}^{k}\mathcal{X}(s_{i},\alpha),$
where $\mathcal{X}(s_{i},\alpha)=\bigcup_{j\in\mathcal{I}_{s_{i}}}\mathcal{X}_{j}(\alpha)$.
Definition 4.8.
Let $\mathfrak{T}$ be a tree graph with the root node $v_{1}$, and $\gamma$ be a path $\gamma=(v_{1},\ldots,v_{k})$ in the tree $\mathfrak{T}$ starting from the root $v_{1}$. The $i^{\textrm{th}}$ partial path of $\gamma$ is defined by
$\gamma_{i}=(v_{1},\ldots,v_{i}),$
where $2\leq i\leq k$, and $\gamma_{1}=v_{1}$.
Then we define a function $\mathcal{M}(\gamma_{i}):[0,1]\times\mathcal{X}\to\mathbb{R}^{N}$, the $i^{\textrm{th}}$ accumulation of votes of $\mathcal{M}$ along with the path $\gamma=(s_{1},\ldots,s_{k})$ in $\mathfrak{T}$ starting from the root $s_{1}$ for $i=1,\ldots,k$. This function maps $\left(\alpha,x\right)\in[0,1]\times\mathcal{X}$ to
(4.3) $\mathcal{M}(\gamma_{i})\rvert_{\alpha,x}:=\mathcal{A}(s_{1},\alpha,x)+\sum_{i=2}^{k}\widetilde{\mathcal{A}}(s_{i},\alpha,x)I_{\mathcal{X}\left(\gamma_{i},\alpha\right)}(x)\textrm{ if }2\leq i,$
where $I_{E}$ denotes the characteristic function of a subset $E\subset\mathcal{X}$ such that
$I_{E}(x)=\begin{cases}1&\textrm{if }x\in E,\\ 0&\textrm{otherwise},\end{cases}$
and $\mathcal{M}(\gamma_{1})\rvert_{\alpha,x}=\mathcal{A}(s_{1},\alpha,x)$.
For a given instance $x\in\mathcal{X}$ and certainty threshold $\alpha\in[0,1]$, we construct a unique path $\gamma$ from the root in the target tree $\mathfrak{T}$ using the function $f:V\times[0,1]\times\mathcal{X}\to[-1,1]$ defined as
(4.4) $f(s,\alpha,x):=\sum_{c_{j}\in\overline{T}_{s}}\mathcal{A}(s,\alpha,x)_{j}-\sum_{c_{j}\notin\overline{T}_{s}}\mathcal{A}(s,\alpha,x)_{j},$
where $V$ is the set of nodes in $\mathfrak{T}$. The path $\gamma$ in $\mathfrak{T}$ from the root $s$ for a given $x$ and $\alpha$ is obtained inductively as follows:
  1. (1)
    Let $s_{1},\ldots,s_{k}$ be the childs of the root $s$. If $f(s_{i},\alpha,x)<0$ for all $i=1,\ldots,k$, then $\gamma=s$. Otherwise,
    $a_{1}:=\min\left\{1\leq i\leq k:f(s_{i},\alpha,x)\geq 0\right\}.$
  2. (2)
    Suppose we have determined $a_{1},\ldots,a_{m}$. If $s_{a_{m}}$ has child nodes $\{s_{1},\ldots,s_{k}\}$, then check if there is a child node $s_{i}$ such that $f(s_{i},\alpha,x)\geq 0$. If no such child node exists, then thew path is $\gamma=(s,s_{a_{1}},\ldots,s_{a_{m}})$ at certainty threshold $\alpha$. Otherwise, define
    $a_{m+1}=\min\left\{1\leq i\leq k:f(s_{i},\alpha,x)\geq 0\right\}.$
Since $\mathfrak{T}$ is a finite graph, this induction terminates in finite steps..
Therefore we have a function $\Gamma:[0,1]\times\mathcal{X}\to\textrm{Path}(\mathfrak{T})$, where $\textrm{Path}(\mathfrak{T})$ is the set of all paths from root nodes, such that $\Gamma(\alpha,x)=\gamma$ is determined by the function $f$ in the Equation 4.4 following the induction steps above. Then define the accumulation of votes of $\mathcal{M}$ at $\alpha$ certainty threshold for $x\in\mathcal{X}$ as the value of $\mathcal{M}\left(\Gamma(\alpha,x)\right)$ at $(\alpha,x)$.
Remark 4.9.
Under the specialization method in 4.2.2, we can construct ‘gating layer’ as in the Mixture of Expert[19] using this voting strategy along the path in the target tree. Let $M$ be a trained model in $\mathcal{M}$, $T=\{c_{1,1},\ldots,c_{1,n_{1}},\ldots,c_{k,n_{k}}\}$, and the logical targets of $M$ is $T_{M}=\left\{t_{1},\ldots,t_{k}\right\}$ where $t_{i}=\{c_{i,1},\ldots,c_{i,n_{i}}\}$ for $i=1,\ldots,k$. Then $M$ serves as the ‘gating’ layer in $\mathcal{M}$ navigating an instance to other trained models that are trained on dataset containing classes exclusively within a logical target of $T_{M}$. See [23] for the experimental results.
4.5.3. Vote using validation history
We introduce the ‘using validation history’ method in prediction to alleviate concerns regarding the optimal accumulation step or certainty threshold. In other words, $\alpha$ and $i$ in the Equation 4.3 are fixed through this method based on the validation dataset used to compute the fuzzy domain of trained models in $\mathcal{M}$. This method operates under the assumption that the given instance follows the same distribution as the validation dataset.
Let $\mathcal{Z}_{\textrm{val}}=\left(\mathcal{X}_{\textrm{val}},\mathcal{Y}_{\textrm{val}}\right)$ denote the validation dataset used to compute the fuzzy domain of each trained model in $\mathcal{M}$, and $\mathfrak{T}$ be the target tree of $\mathcal{M}$. Let $V$ denote the set of nodes in $\mathfrak{T}$. For $v\in V$ and $\alpha\in[0,1]$, we define $D(v,\alpha)\subset\mathcal{X}_{\textrm{val}}$ as the set of points $x\in\mathcal{X}_{\textrm{val}}$ such that the path $\Gamma(\alpha,x)$ contains $v$. For any path $\gamma\in\textrm{Path}(\mathfrak{T})$, we have $\mathcal{M}(\gamma)\in\mathbb{R}^{N}$ and therefore the prediction made by $\mathcal{M}(\gamma)$ is $c_{\arg\max\mathcal{M}(\gamma)}\in T$. Let $P(\gamma)$ denote the prediction made by $\mathcal{M}(\gamma)$. Then we define the expected accuracy $r_{(v,\alpha)}$ at $(v,\alpha)\in V\times[0,1]$ as
|$\frac{\left|\left\{x\in D(v,\alpha):\left(P(\gamma_{i}),y\right)\in\mathcal{Z}\textrm{ where }\Gamma(\alpha,x)=\gamma=(s_{1},\ldots,s_{k})\textrm{ and }v=s_{i}\right\}\right|}{\left|D(v,\alpha)\right|},$
which represents the accuracy of $\mathcal{M}$ over the sample that follows a path in the target tree $\mathfrak{T}$ targeted to $v$ at certainty threshold $\alpha$. In practice, the validation dataset $\mathcal{Z}_{\textrm{val}}$ is finite set, and therefore the range of $r_{(v,\alpha)}$ is finite. Let the validation history be the set of all expected accuracies. For a given instance $x\in\mathcal{X}$, define a set $H_{x}$ of expected accuracies from the validation history as
$H_{x}:=\{r_{s,\alpha}\mid s\textrm{ is a node contained in the path }\Gamma(\alpha,x),\alpha\in[0,1]\}.$
Let $s^{*}$ and $\alpha_{*}$ be a node and a certainty threshold at which the expected accuracy attains its maximum in $H_{x}$. Then $s^{*}\in\Gamma(\alpha_{*},x)$, where $\Gamma(\alpha_{*},x)$ is denoted by $(v_{1},\ldots,v_{k})$ with $v_{i}=s^{*}$, and the vote using the validation history is defined by $\mathcal{M}\left(\gamma_{i}\right)$ where $\gamma=\Gamma(\alpha_{*},x)=(v_{1},\ldots,v_{k})$.
Remark 4.10.
Let a system of trained models $\mathcal{M}$ and validation dataset $\mathcal{Z}_{\textrm{val}}$ be given. The combining rule for each comprising trained model using validation history serves as the role of $\rho$ in the Definition 3.13, and $P:\mathcal{X}\times T\to[0,1]$ is defined by the vote of $\mathcal{M}$ using the validation history.
Appendix A Pseudo-Algorithms
In this appendix we summarize implementations in section 4 with pseudo-algorithms. We follow the frequently used symbols in the Table 1.
Algorithm 1 Finding Fuzzy Domain
1:A trained model $M$, Dataset $\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$,
2:fuzzy domain =$\mathcal{F}=\{F_{\alpha_{0}},\ldots,F_{\alpha_{n}}\}$
3:Set $P={{\left(\arg\max M(x)\right)}}_{x\in\mathcal{X}}$triangleright predictions
4:Set $C={(\max M(x))}_{x\in\mathcal{X}}$triangleright Certainty
5:Set $T=\{t_{1},\ldots,t_{k}\}$triangleright The logical targets
6:Set $\alpha=0$.
7:repeat
8:     $F_{\alpha}\leftarrow\left(\Phi_{\mathcal{X}}(\alpha),\frac{\left|\mathcal{X}_{M}(\alpha)\right|}{\left|\mathcal{X}\right|},\left(\phi_{\mathcal{X},t_{1}}(\alpha),\ldots,\phi_{\mathcal{X},t_{k}}(\alpha)\right)\right)$
9:     Increase $\alpha$ slightly
10:until $\left|\mathcal{X}_{M}(\alpha)\right|$ is sufficiently small or $\alpha\geq 1$
11:return ${\{F_{\alpha}\}}_{\alpha\in\{0=\alpha_{0},\alpha_{1},\ldots,\alpha_{n}\}}$
Algorithm 2 Prediction in a Flat Target
1:Thresholds $A=\{0=\alpha_{0},\ldots,\alpha_{n}\}$, Target classes $T=\{c_{1},\ldots,c_{N}\}$, trained models $M_{1},\ldots,M_{k}$, Dataset $\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$ Assumption : Each target classes $T_{i}=\{t_{i,1},\ldots,t_{i,N_{i}}\}$ of $M_{i}$ has the same flattening $T$.
$\mathcal{I}\leftarrow$ a set of labels $k$ of $M_{1},\ldots,M_{k}$
3:for $i$ runs over labels $\mathcal{I}$ do
4:     Set $a_{c,i}$ be $1/|t|$, the weight to class $c$ where $c\in t\in T_{i}$
5:     Set $\widehat{T}_{i}$ be the categorized matrix of dimension $|T_{i}|\times|T|$ converted from $T_{i}$ using one-hot encoding, whose $(u,v)-$ component is 00 if $c_{v}\notin t_{u}$, and $a_{c_{v},i}$ if $c_{v}\in t_{u}$.
6:end for
7:for $\alpha$ runs over thresholds $A$ do
8:     for $i$ runs over labels $\mathcal{I}$ do
9:         $\Phi_{\mathcal{X},i}(\alpha),\mathcal{X}_{i}(\alpha),\left(\phi_{\mathcal{X},i,t_{i,j}}(\alpha)\right)_{j=1,\ldots,N_{i}}\leftarrow\textsc{Get Fuzzy Domain}(M_{i},\mathcal{Z})$
10:         for $x\in\mathcal{X}$ do
11:              $P_{i}(x)=$ the prediction of $M_{i}$ for $x$
12:              if  $P_{i}(x)=t_{v}$ and $x_{u}\in\mathcal{X}_{i}(\alpha)$ then Set $\left[\mathfrak{P}_{i}(\alpha)\right]_{u,v}=\phi_{\mathcal{X},i,t_{i,v}}(\alpha)$
13:              else if $P_{i}(x)=t_{v}$ and $x_{u}\in\mathcal{X}\setminus\mathcal{X}_{i}(\alpha)$ then Set $\left[\mathfrak{P}_{i}(\alpha)\right]_{u,v}=\phi_{\mathcal{X},i,t_{i,v}}(0)$
14:              else$\left[\mathfrak{P}(\alpha)\right]_{u,v}=0$
15:              end if
16:         end for
17:         Set a matrix $\mathfrak{P}_{i}(\alpha)$ whose $(u,v)-$ component is $\left[\mathfrak{P}_{i}(\alpha)\right]_{u,v}$ where $v=1,\ldots,N_{i}$.
18:     end for
[…]
21:end for
22:$\left(\text{Votes}_{\alpha}(x):\alpha\in A,x\in\mathcal{X}\right)$, $\left(\text{CertainPart}_{\alpha}:\alpha\in A\right)$
Algorithm 3 Specialization
1:A trained model $M$ on dataset $\mathcal{Z}=\left(\mathcal{X},\mathcal{Y}\right)$, Target of classes $\{c_{1},\ldots,c_{N}\}$, New targets $T=\{t_{1},\ldots,t_{m}\}$
2:Initialize an array of tuples $\mathcal{Z}_{\textrm{new}}$
3:for $t_{i}$ runs over new targets $T=\{t_{1},\ldots,t_{m}\}$ do
4:     Set $\mathcal{X}_{t_{i}}=\underset{c\in t_{i}}{\bigcup}\{x\in\mathcal{X}|(x,c)\in\mathcal{Z}\}$
5:     Add $\left(x,t_{i}\right)_{x\in\mathcal{X}_{t_{i}}}$ to $\mathcal{Z}_{\textrm{new}}$
6:end for
7:$base\leftarrow$ new model that takes the input of $M$ outputs the result of the second-to-last layer.
8:$M_{new}\leftarrow base$ with a layer having $m$ number of outputs and Softmax layer
9:Train $M_{\textrm{new}}$ on the $\mathcal{Z}_{\textrm{new}}$
10:Specialized model $M_{\textrm{new}}$ with target $T$